### v7

# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

Abstract
[Objective] This study focuses on addressing the current lack of a unified formal theoretical framework in machine learning, as well as the deficiencies in interpretability and ethical safety assurance.
[Methods] A formal information model is first constructed, utilizing sets of well-formed formulas to explicitly define the ontological states and carrier mappings of typical components in machine learning. Learnable and processable predicates, along with learning and processing functions, are introduced to analyze the logical deduction and constraint rules of the causal chains within models.
[Results] A meta-framework for machine learning theory (MLT-MF) is established. Based on this framework, universal definitions for model interpretability and ethical safety are proposed. Furthermore, three key theorems are proved: the equivalence of model interpretability and information recoverability, the assurance of ethical safety, and the estimation of generalization error.
[Limitations] The current framework assumes ideal conditions with noiseless information-enabling mappings and primarily targets model learning and processing logic in static scenarios. It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems.
[Conclusions] This work overcomes the limitations of fragmented research and provides a unified theoretical foundation for systematically addressing the critical challenges currently faced in machine learning.
Keywords: machine learning; objective information theory; formal logic systems; interpretability; ethical and safety assurance; generalization error
1 Introduction
Although artificial intelligence is currently the most dynamic field within information science and even the broader domain of science and technology, existing research still faces several significant shortcomings:
First, there is a lack of a formalized framework that covers the entire lifecycle of machine learning [1, 2, 3, 4, 5]. Current theoretical studies in machine learning tend to focus on isolated components: optimization theory emphasizes training dynamics, generalization analysis concerns test performance, and strategies for initialization and deployment are often treated as independent stages. This fragmentation makes it impossible to characterize the coherent evolution of knowledge from prior injection, through data-driven updates, to task generalization. For example, the PAC-Bayes framework connects priors with generalization error but neglects the temporal causality of the training process; meta-learning involves multi-stage adaptation but fails to establish a symbolic model for state transitions. The absence of lifecycle modeling makes systematic analysis of the core question——“how parameters carry task knowledge from initialization through training”—difficult for model designers.
Second, there is a lack of explicit expression for the causal chain of information flow. Traditional approaches often treat the transformation from data to model as a black box, reducing the parameter update driven by loss function gradients to a purely mathematical optimization process and overlooking the causal dependencies among prior knowledge, data distribution, and model architecture. While the field of causal inference attempts to characterize causal relationships among variables, its application in machine learning is mostly limited to feature-label association analysis and fails to construct a complete causal chain from “prior constraints → initialization → gradient propagation → generalization capability.” For instance, adversarial training enhances robustness by maximizing loss, but theoretical analyses rarely explicitly link initial parameter constraints with the causal mechanisms underlying final adversarial robustness, making strategy design reliant on empirical trial-and-error.
Third, the mathematical foundations of interpretability tools are weak [11, 12, 13, 14, 15]. Most current interpretability methods use local approximation or game-theoretic approaches to assign feature importance, which, while practically useful, lack rigorous mathematical axiomatization. For example, SHAP values compute marginal contributions in cooperative games but do not formalize their mapping to the information carrier states of models, potentially violating causality in interpretations. In addition, heuristic algorithms that infer data-feature dependencies from parameters lack mathematical guarantees for “information recoverability.” This theoretical weakness makes interpretability results susceptible to adversarial attacks and limits their applicability in high-stakes scenarios such as medicine or law.
Information science, as a core discipline connecting data, knowledge, and systems, offers a new perspective for the theoretical integration of machine learning[16, 17] . Objective Information Theory (OIT) establishes a sextuple model of information through axiomatic methods, representing the intrinsic meaning of information via ontological states, reflecting its objective expression through carrier states, and explaining information realization through enabling mappings that possess both mathematical surjectivity and physical causality. This provides a universally applicable foundational model for understanding information in the objective world, especially within computer and information systems[18, 19, 20, 21, 22, 23].
On this basis, this paper rigorously defines higher-order logical systems using recursive methods and establishes a formalized expression of states via the interpretation of sets of well-formed formulas, thus fully clarifying the mathematical essence of state representation in the definition of information. This fills the last theoretical gap in OIT and provides more powerful tools for applying OIT to analyze various problems in information science. Subsequently, the theories and methods of OIT are applied comprehensively—from prior knowledge encoding and data-driven training to model generalization applications—unifying core concepts such as prior distribution, data distribution, and parameter updates into a formalized state representation for the first time. This establishes a mathematical logic system that spans the entire machine learning process. Based on this, formal logical deduction and constraint rules for the causal chains in typical stages such as formal modeling and realization of training information, formal modeling and realization of model learning, and information processing and learning optimization in models are demonstrated, forming the Meta-Framework for Machine Learning Theory (MLT-MF).
Finally, MLT-MF is applied to several hot topics in machine learning, establishing universal definitions for model interpretability and ethical safety, and proving three important theorems: the equivalence of model interpretability and information recoverability, ethical safety assurance, and generalization error estimation. This provides a powerful tool based on a unified formal logical framework for both theoretical analysis and engineering practice in machine learning.
2 Postulates and Formal Definition of Information
Undoubtedly, as a key component of artificial intelligence, modern machine learning is a product of the rapid development and high integration of information technologies such as computers, the Internet, big data, and cloud computing. Machine learning systems are, without question, an important branch of information systems. Starting from the basic concept of information can provide a unified theoretical foundation for research in machine learning and artificial intelligence.
2.1 Postulates of Information
Although humanity has entered the information age, the theoretical foundations of information science have long been dispersed across disciplines such as information theory, cybernetics, signal processing, statistical learning, and computational theory, each developing independently. There remains a lack of consensus on the concept of information, as well as a unified axiomatic basis and theoretical system. This is also a major reason why artificial intelligence often lacks interpretability and predictability. In response,[22] proposes four Postulates of information:
Postulate 1 (Binary Attribute) Every piece of information comprises two components: the ontology, which represents the intrinsic essence of the information, and the carrier, which embodies the objective form of the information.
Postulate 2 (Existence Duration) The content and form of information have their own respective temporal existences.
Postulate 3 (State Representation) The states of the ontology and the carrier of information have their own respective representations.
Postulate 4 (Enabling Mapping) The state of the ontology can be mapped onto the state of the carrier via an enabling mapping.
These four postulates are concise, align with common conceptions of information, and are compatible with the basic concepts of major branches of information science. Their scientific significance is reflected in:
  • Unification: Providing a common foundation for long-dispersed theories in information science;
  • Axiomatization: Allowing the definition and reasoning of information to be axiomatized and formalized, akin to mathematics and physics;
  • Universality: Facilitating application to cutting-edge fields such as the Internet, big data, artificial intelligence, cognitive science, and cyber-physical systems.
The concept of "enabling mapping" is novel. It requires not only that a surjective mapping from the ontological state to the carrier state be established mathematically, but also that the existence of the ontological state renders the carrier state an objective reality, highlighting the dual character of information as both a mathematical mapping and a physical realization[24, 25, 26].
2.2 Formal Representation of States
The postulates of information emphasize that the state of the ontology enables a mapping to the state of the carrier. Therefore, establishing a formal representation of both ontological and carrier states is particularly important.
Definition 1 (Higher-Order Formal Logical System): Suppose a formal system $\mathcal{L}$ is described using the following symbol set:
Variables $x_{1},x_{2},\ldots$;
Individual constants $a_{1},a_{2},\ldots$;
Functions $f_{1}^{1},f_{2}^{1},\ldots,f_{1}^{2},f_{2}^{2},\ldots,f_{1}^{3},f_{2}^{3},\ldots$;
Parentheses (,)(,)( , );
Predicates $A_{1}^{1},A_{2}^{1},\ldots,A_{1}^{2},A_{2}^{2},\ldots,A_{1}^{3},A_{2}^{3},\ldots$;
Logical connectives "not" $\sim$, "implication" $\rightarrow$;
Quantifier "for all" $\forall$;
A term in $\mathcal{L}$ is defined as follows:
(1) Variables and individual constants are terms;
(2) If $f_{i}^{n}(n>0,i>0)$ is a function in $\mathcal{L}$, and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ that do not contain $f_{i}^{n}$ itself, then $f_{i}^{n}(t_{1},\ldots,t_{n})$ is also a term in $\mathcal{L}$;
(3) If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$, and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ that do not contain $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1},\ldots,t_{n})$ is a term in $\mathcal{L}$.
(4) The set of all terms in $\mathcal{L}$ is generated by (1), (2), and (3).
Based on this, we can define atomic formulas in $\mathcal{L}$:
If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$ and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ not containing $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1},\ldots,t_{n})$ is an atomic formula in $\mathcal{L}$.
We then define well-formed formulas (WFFs) in $\mathcal{L}$:
(5) Every atomic formula in $\mathcal{L}$ is a WFF;
(6) If $\mathcal{A}$ and $\mathcal{B}$ are WFFs in $\mathcal{L}$, then $\sim\mathcal{A}$, $\mathcal{A}\rightarrow\mathcal{B}$, $(\forall x_{i})\mathcal{A}$, $(\forall f_{i}^{n})\mathcal{A}$, and $(\forall A_{i}^{n})\mathcal{A}$ are all WFFs in $\mathcal{L}$, where $x_{i}$, $f_{i}^{n}$, $A_{i}^{n}$ are variables, functions, and predicates in $\mathcal{L}$, respectively, and no predicate or function symbol is applied, either directly or indirectly, to itself.
(7) The set of all WFFs in $\mathcal{L}$ is generated by (5) and (6).
The higher-order formal logical system $\mathcal{L}$ defined here, based on the flexibility of predicate logic, can describe various relationships and constraints among objects in a unified way. Through higher-order quantification support, the nested definition of complex data and behaviors, and it explicitly prohibits anypredicate or function symbol from applying to itself—directly or indirectly—to avoid self-referential paradoxes.
For simplicity, $\mathcal{L}$ does not define connectives such as "and" $\land$, "or" $\lor$, "if and only if" $\leftrightarrow$, or the quantifier "exists" $\exists$, and other commonly used logical symbols. However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of symbols and can make many formulas more concise[27]. Therefore, they are introduced as defined symbols in subsequent discussions.
Definition 2 (Interpretation of a Formal System): An interpretation $E$ of formal system $\mathcal{L}$ over a domain $D_{E}$ consists of a non-empty set $D_{E}$, a specified set of elements $(\widehat{a}_{1},\widehat{a}_{2},\ldots)$ over $D_{E}$, a set of functions $(\widehat{f}_{i}^{n},n>0,i>0)$ over $D_{E}$, and a set of relations $(\widehat{A}_{i}^{n},n>0,i>0)$ over $D_{E}$ and the specified elements.
According to this definition, the variables $x_{1},x_{2},\ldots$ in $\mathcal{L}$ are interpreted as objects. The set $D_{E}$ is the domain over which variables take their values. The set of specfied elements corresponds to the special objects denoted by the individual constants in $\mathcal{L}$. Functions and relations provide concrete interpretations for the function and predicate symbols in $\mathcal{L}$.
Definition 2 assigns concrete values and meaning to a set of formulas in a formal system $\mathcal{L}$ through interpretation, thereby providing a fundamental approach for formally describing the definitional, attributive, structural, relational, and evolutionary state properties of objects as both the subjective and objective worlds.
Postulate 5 (Expression of Object State):The state set of an object collection over a specified time set is constituted by its attributes—such as properties, structures, values, and relations—as well as the evolution rules and logical combinations among these attributes, and satisfies the following conditions:
(1) Definitional Invariance: The definitions of the object set and the time set remain invariant over the entire domain.
(2) Expressivity of Attributes: All attributes of the object set—including properties, structures, values, and relations—over the domain can be expressed via functions and predicates within a formal system conforming to Definition 1.
(3) Expressivity of Rules: The evolution rules governing various attributes of the object set can be expressed, without implicit self-reference, via temporal variables and application connectives in a formal system as specified in Definition 1.
(4) Logical Consistency: There exists no logical contradiction among the various attributes and evolution rules of the object set, nor among any possible logical conjunctions thereof.
Theorem 1 (Formal Expression of States): Let $S(X,T)$ be the state set of an object collection $X$ over a specified time set $T$, satisfying Postulate 5. Then $S(X,T)$ can always be represented as the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$, and is logically consistent.
Proof: Since $S(X,T)$ satisfies Postulate 5, we have:
  • By (1), the elements of $X\times T$ are well-defined, so $S(X,T)$ has a definite domain.
  • By (2), all attributes of $X$ over $T$ in $S(X,T)$ can be expressed by functions and predicates in a formal system $\mathcal{L}$ as specified in Definition 1.
  • By (3), suppose $A_{1}(x_{1},t_{1},f_{1}(x_{1},t_{1}))$ and $A_{2}(x_{2},t_{2},f_{2}(x_{2},t_{2}))$ are two attributes in $S(X,T)$ pertaining to objects $x_{1},x_{2}\in X$, times $t_{1},t_{2}\in T(t_{1}<t_{2})$, and functions $f_{1}(x_{1},t_{1}),f_{2}(x_{1},t_{1})$. If every $A_{1}(x_{1},t_{1},f_{1}(x_{1},t_{1}))$ evolves from $t_{1}$ to $t_{2}$ according to some rule into $A_{2}(x_{2},t_{2},f_{2}(x_{2},t_{2}))$,
    then $(\forall A_{1})(\forall f_{1})(\forall x_{1})(A_{1}(x_{1},t_{1},f_{1}(x_{1},t_{1}))\rightarrow\exists A_{2}\exists f_{2}\exists x_{2}(A_{2}(x_{2},t_{2},f_{2}(x_{2},t_{2}))))$ is a well-formed formula in $\mathcal{L}$ that, as an interpretation over $x_{1},x_{2},t_{1},t_{2}$, expresses the corresponding rule in $S(X,T)$. Other rules in $S(X,T)$ can likewise be expressed through corresponding well-formed formulas in $\mathcal{L}$ interpreted over $X\times T$.
  • Let the set of such well-formed formulas in $\mathcal{L}$ representing $S(X,T)$ be defined as its interpretation ${E}$ over the domain $X\times T$.
  • By (4), no logical contradiction exists among any attributes or rules in $S(X,T)$, nor among any logical conjunctions thereof. Thus, as a representation of $S(X,T)$, both ${E}$ and its logical closure are logically consistent.
In summary, $S(X,T)$ is indeed an interpretation ${E}$ of a set of well-formed formulas in $\mathcal{L}$ over the domain $X\times T$, and both itself and its logical closure are logically consistent. The theorem is thus proved.
Theorem 1 indicates that, by employing the higher-order formal system $\mathcal{L}$, it is possible to describe the states of virtually all things in the world. Both theoretical and empirical analyses [28] further demonstrate that $\mathcal{L}$ is broadly applicable to the expression of state in disciplines such as mathematics, economics, ecology, computer and information systems, and natural language. This provides a unified state modeling theory for interdisciplinary research, which is of significant importance to fields such as artificial intelligence, knowledge engineering, and complex systems.
2.3 Finite Automaton States
Using the formal expression of states, we can study in depth the special and important object of finite automata.
Theorem 2 (State Representation of Finite Automata): For any finite automaton $M$, one can describe its input, output, and state-transition behavior over a related time set $T$ by the state set $S(M,T)$.
Proof: Let $M=(Q,R,U,\delta,\lambda)$ be a finite automaton, and $T=\{t_{i}|i=1,\ldots,n_{T}\}$ the (strictly increasing) time points at which $M$ performs input, output, or state-transition actions. By finite automaton theory [29], $Q=\{q_{t_{i}}|t_{i}=1,\ldots,n_{T}-1\}$, $R=\{r_{t_{i}}|t_{i}=2,\ldots,n_{r}\}$, and $U=\{u_{t_{i}}|t_{i}=1,\ldots,n_{u}\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
$\delta:U\times Q\to U$ is the next-state function;
$\lambda:U\times Q\to R$ is the output function.
For each state $u_{t_{i}}\in U$, define the state predicate $State^{3}$ by
$\varphi_{U}(t_{i})=State^{3}(M,t_{i},u_{t_{i}})$ (2.1)
to assert that at time $t_{i}$, $M$ is in state $u_{t_{i}}$.
For each input $q_{t_{i}}\in Q$, define the input predicate $Input^{3}$ by
$\varphi_{I}(t_{i})=Input^{3}(M,t_{i},q_{t_{i}})$ (2.2)
to assert that at time $t_{i}$, $M$ receives input $q_{t_{i}}$.
For each output $r_{t_{i}}\in R$, define the output predicate $Output^{3}$ by
$\varphi_{O}(t_{i})=Output^{3}(M,t_{i},r_{t_{i}})$ (2.3)
to assert that at time $t_{i}$, $M$ produces output $r_{t_{i}}$.
The transition function $\delta$ is captured by the well-formed formula
$\varphi_{\delta}(t_{i})=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\delta(u_{t_{i}},q_{t_{i}})\land Eq^{2}(u_{t_{i+1}},\delta(u_{t_{i}},q_{t_{i}}))\land\varphi_{U}(t_{i+1})$ (2.4)
where the binary predicate $Eq^{2}(x,y)$ asserts $x=y$.
Similarly, the output function $\lambda$ is given by
$\varphi_{\lambda}(t_{i})=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\lambda(u_{t_{i}},q_{t_{i}})\land Eq^{2}(r_{t_{i+1}},\lambda(u_{t_{i}},q_{t_{i}}))\land$ (2.5)
$IsElementof^{2}(r_{t_{i+1}},R)\land\varphi_{O}(t_{i+1})$
where $IsElementof^{2}(x,X)$ asserts $x$ is an element of $X$.
Thus the state set
$S(M,T)=Cn(\{\varphi_{U}(t_{i})\}\cup\{\varphi_{I}(t_{i})\}\cup\{\varphi_{O}(t_{i})\}\cup\{\varphi_{\delta}(t_{i})\}\cup\{\varphi_{\lambda}(t_{i})\})$ (2.6)
describes $M$’s inputs, outputs, and state-transitions over $T$. Here $Cn(X)$ denotes the logical closure of the set of well-formed formulas $X$. This completes the proof.
It follows that every finite automaton can be formalized by the state set given in Theorem 2, which moreover exhibits atuomatic transtion behavior. Hence we introduce :
Definition 3 (Finite Automaton State): If a state set $S(X,T)$ expresses a finite automaton, then $S(X,T)$ is called a finite automaton state.
All practical information systems, including computing systems, are finite automata in the mathematical sense. Therefore, the concept of finite automaton state is of great significance in the study of machine learning.
2.4 Definition of Information
As shown in[22], the necessary and sufficient conditions for satisfying Postulates 1-4 are the existence of the following non-empty sets: an ontological set $o$, a set of times for the occurrence of the ontology $T_{h}$, a set of ontological states $S_{o}(o,T_{h})$, a set of objective carriers $c$, a set of times for the reflection of the carrier $T_{m}$, and a set of carrier reflection states $S_{c}(c,T_{m})$, such that $S_{o}(o,T_{h})$ enables a mapping to $S_{c}(c,T_{m})$. Denote this mapping as $I$, or more simply, $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, which is called the sextuple model of information.
This leads to the general definition of information:
Definition 4 (Mathematical Definition of Information): Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain. The ontology $o\in\mathbb{O}\cup\mathbb{S}$, time of occurrence $T_{h}\subseteq\mathbb{T}$, ontological state set $S_{o}(o,T_{h})$ satisfying Postulate 5, carrier $c\subseteq\mathbb{O}$ , reflection time $T_{m}\subseteq\mathbb{T}$, and carrier reflection set $S_{c}(c,T_{m})$ satisfying Postulate 5 are all non-empty sets. Then, information $I$ is the enabling mapping from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$, that is,
$I:S_{o}(o,T_{h})\rightrightarrows S_{c}(c,T_{m})$ (2.7)
or
$I\left(S_{o}(o,T_{h})\right)\underline{\triangleright}S_{c}(c,T_{m})$ (2.8)
The set of all information $I$ is called the information space, denoted by $\mathfrak{I}$, and is one of the three major constituents of the objective world.
[…]
This shows that the carrier state set $S_{c}(c,T_{m})$ is always induced by the ontological state set $S_{o}(o,T_{h})$, and thus the enabling mapping $I$ must be surjective. For simplicity, unless otherwise noted, subsequent discussions assume that $I$ is a single-valued mapping.
Therefore, for any $\psi\in S_{c}(c,T_{m})$, there exists at least one $\varphi\in S_{o}(o,T_{h})$ such that $I(\varphi)\underline{\triangleright}\psi$. In this case, we write $I^{-1}(\psi)\underline{\triangleleft}\varphi$ or $\varphi\underline{\triangleright}I^{-1}(\psi)$ to preserve the consistency of the enabling direction.
It is important to note that both $S_{o}(o,T_{h})$ and $S_{c}(c,T_{m})$ are state representations that satisfy Postulate 5. According to Theorem 1, each can be interpreted as a set of well-formed formulas within some formal system. However, the corresponding formal systems for the ontology and the carrier need not be identical. This distinction reflects one of the essential characteristics of information: its ability to establish connections between disparate entities in the world.
Definition 4 does not require that the ontological states of information be mapped one-to-one onto the carrier states. Of course, in practical scenarios, one-to-one mappings often do exist, which can simplify the analysis considerably.
Definition 5 (Recoverable Information): Let $I=(o,T_{h},S_{o},c,T_{m},S_{c})$ still denote a bijection from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$, i.e., $I$ is invertible. Thus, there exists an inverse mapping $I^{-1}$ such that $I^{-1}(S_{c}(c,T_{m}))\underline{\triangleleft}S_{o}(o,T_{h})$. In this case, we say that information $I$ is recoverable, and $S_{o}(o,T_{h})$ is the recoverable state set of information $I$.
In particular, owing to the importance of information in information systems, we propose:
Definition 6 (Information in an Information System): Let $I=(o,T_{h},S_{o},c,T_{m},S_{c})$ be a recoverable information. If $S_{o}(o,T_{h})$ is a finite automaton state, then $c$ is called an information system, and $I$ is called information in the information system $c$.
For recoverable information, we also use $I^{-1}(S_{c}(c,T_{m}))\underline{\triangleleft}S_{o}(o,T_{h})$ or $I^{-1}:S_{c}(c,T_{m})\leftleftarrows S_{o}(o,T_{h})$ to maintain consistency in the direction of enabling. Since, as recoverable information, the carrier losslessly carries the complete state of the ontology, this naturally represents an ideal scenario in information studies. If $I$ is not a bijection, we can always partition elements of $S_{o}(o,T_{h})$ according to the same image equivalence classes, and then establish an enabling mapping from these equivalence classes to $S_{c}(c,T_{m})$ by means of equivalence class reduction, thereby ensuring recoverability. Specifically:
Proposition 1 (Recoverable Reduction of Information): For $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, let $R$ be an equivalence relation on $S_{o}(o,T_{h})$ defined by $xRy\leftrightarrow I(x)=I(y)$. Then, there exists a unique enabling mapping $\hat{I}:S_{o}(o,T_{h})/R\rightrightarrows S_{c}(c,T_{m})$, and $\hat{I}$ is a bijection, i.e., $\hat{I}$ is recoverable. In this case, $\hat{I}$ is called the recoverable reduction of information $I$.
When $I$ is already recoverable, $S_{o}(o,T_{h})/R=S_{o}(o,T_{h})$, and $\hat{I}=I$. Recoverable information and the recoverable reduction of information can provide significant assistance in simplifying problems.
Reference [24]is generated as a foundational work in information theory. The paper proposed that "$H=-\sum p_{i}\log p_{i}$ plays a central role as the measure of information in information theory." The discussion in this paper includes both the source and channel of information. The source is the origin of the information, and it generates events $e_{1},e_{2},\ldots,e_{n}$ with probabilities $p_{1},p_{2},\ldots,p_{n}$, respectively. The channel $I_{c}$ is the medium through which information is transmitted. After an event $e_{i}$ occurs at time $T_{h}$ in the source $I_{s}$, the channel $I_{c}$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$(where $01\leq i\leq n,\tau>$ 0).We define the ternary predicate $IsEvent^{3}(I_{s},T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $I_{s}$ at time $T_{h}$, and the ternary predicate $IsCode^{3}(I_{c},T_{m},c_{i})$ to represent that the channel $I_{c}$ received and transmitted code $c_{i}$ in time interval $T_{m}$. Then,
$S_{I_{s}}(I_{s},T_{h})=IsEvent^{3}(I_{s},T_{h},e_{i})$ (2.9)
and
$S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))=IsCode^{3}(I_{c},(T_{h},T_{h}+\tau),c_{i})$ (2.10)
are the states of the source $I_{s}$ at time $T_{b}$ and the state of the channel $I_{c}$ in the time interval $(T_{h},T_{h}+\tau)$, respectively (where $1\leq i\leq n)$. Furthermore,
$I:S_{I_{s}}(I_{s},T_{h})\rightrightarrows S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$ (2.11)
represents the functional enabling mapping from the source state $S_{I_{s}}(I_{s},T_{h})$ to the channel state $S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$.From equation (2.10), we can recover the event $e_{i}$ in equation (2.9) using the code $c_{i}$, thus $I$ is still recoverable information. Therefore, although Shannon’s information theory investigates the various probability distributions of $S_{I_{s}}(I_{s},T_{h})$ and the complex implementation mechanisms of $S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$, it forms a rich and extensive theoretical system. The OIT definition of information can encompass Shannon’s most original model of information, indicating that it has a broader and more universal application range.
2.5 Noisy Information
Many researchers question the significance of recoverable information, as in the real world, information is often not perfectly recoverable. This is because the reflection of information is often influenced by noise or other external factors. Therefore, the following concept is introduced:
Definition 7 (Noisy Information): For information $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, let $\widetilde{S_{o}}(o,T_{h})$ be another state of the same carrier $c$ at time $T_{m}$ and $\widetilde{S_{c}}(c,T_{m})$ be another state of the same carrier $c$ at time $T_{m}$,
$\widetilde{S_{o}}(o,T_{h})=S_{o}(o,T_{h})\backslash S_{o}^{-}(o,T_{h})\cup S_{o}^{+}(o,T_{h})$ (2.12)
where the state $S^{-}_{o}(o,T_{h})$ is the loss state of $S_{o}(o,T_{h})$, satisfying $S^{-}_{o}(o,T_{h})\subseteq$ $S_{o}(o,T_{h})$ and not being enabled to map to $\widetilde{S_{c}}(c,T_{m})$, and the state $S^{+}_{o}(o,T_{h})$ is the superposed state of $S_{o}(o,T_{h})$, satisfying $S_{o}(o,T_{h})\cap S_{o}^{+}(o,T_{h})=\emptyset$ and being enabled to map to $\widetilde{S_{c}}(c,T_{m})$. The enabling mapping is then defined as:
$\tilde{I}:\widetilde{S_{o}}(o,T_{h})\rightrightarrows\widetilde{S_{c}}(c,T_{m})$ (2.13)
which is the noisy information of $I$, while $I$ is its own noiseless information. Thus, the noisy information $\tilde{I}$ is also information. It has the same ontology, time of occurrence, carrier, and reflection time as $I$, but the states of the entity and carrier may overlap but are not completely the same. Further analysis of equation (2.12) shows that information $I$ and $\tilde{I}$ are completely asymmetric, and hence:
Proposition 2 (Symmetry of Noisy Information): If $\tilde{I}$ is the noisy information of $I$, then $I$ is also the noisy information of $\tilde{I}$.
In fact, for both information $I$ and its noisy information $\tilde{I}$, it is easy to verify its recoverability or recoverable reducibility from the perspective of enabling mapping. However, in real scenarios, it is often difficult to obtain noise-free information $I$, and one can only study using a noisy version $\tilde{I}$. Moreover, the recoverable state of $\tilde{I}$ is not the same as that of $I$, which is the fundamental reason for the skepticism about the recoverability of information. But under the condition where the noiseless information $I$cannot be directly obtained, we can approximate the ideal state of $I$ using $\tilde{I}$ along with the relationships between $S_{o}(o,T_{h}),S^{-}_{o}(o,T_{h})$, and $S_{o}^{+}(o,T_{h})$. This is the practical significance of introducing the concept of noisy information.
Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenario[24]. In this case, Shannonassumes that the signal generated by the source is $S$, the channel noise is $N$, and the channel output is:
$E=f(S,N)$ (2.14)
where $f$ is the distortion function of the input for the transmitted signal. According to OIT’s view, the signal $S$ here is the ontological state of information $I$, $S_{I_{s}}(I_{s},T_{h})$. The distortion function $f(S,N)$ for signal $S$ and noise $N$ is equivalent to removing a state $S_{I_{s}}^{-}(I_{s},T_{h})$ from the source state $S_{I_{s}}(I_{s},T_{h})$ and adding another state $S_{I_{s}}^{+}(I_{s},T_{h})$. The output $E$ is then the channel’s output state $\widetilde{S_{I_{c}}}(I_{c},(T_{h},T_{h}+\tau))$, where $(T_{h},T_{h}+\tau)$ is the time interval of the channel output. Thus,
$\tilde{I}:\widetilde{S_{I_{s}}}(I_{s},T_{h})\rightrightarrows\widetilde{S_{I_{c}}}(I_{c},(T_{h},T_{h}+\tau))$ (2.15)
is the noisy information of
$I:S_{I_{s}}(I_{s},T_{h})\rightrightarrows S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$ (2.16)
where $\widetilde{S_{I_{c}}}(I_{s},T_{h})=S_{I_{s}}(I_{s},T_{h})\backslash S_{I_{s}}^{-}(I_{s},T_{h})\cup S_{I_{s}}^{+}(I_{s},T_{h})$.
In this way, we use the definition of noisy information to explain Shannon’s model of noisy channels in[24], demonstrating that OIT is entirely applicable to the complex scenarios of information research under noisy conditions.
3 Meta-Framework for Machine Learning Theory (MLT-MF)
The core of machine learning is the model, which is a typical finite automaton and also an information system. Therefore, every learning process and the subsequent testing or application process rely on the information within it, which can generally be summarized and divided into four main steps of information realization and operation.
Step 1: The implementation of training information $It$, in which the ontological state of $It$ is enabled and mapped onto the model interface as its carrier;
Step 2: The original information $Io$ within the model—endowed with learning capability—is fused with the training information $It$, and optimized to yield the updated information $Iu$, which possesses both the ability to answer specific tasks and the capacity for further learning;
Step 3: The implementation of user input information $Iq$, where the ontological state of $Iq$ is likewise enabled and mapped onto the model interface as its carrier;
Step 4: The updated information $Iu$ within the model processes the input $Iq$, outputs the response information $Ir$ according to user requirements, and simultaneously undergoes further optimization through learning, resulting in $Iv$.
3.1 Formal Modeling and Realization of Training Information
According to the general formal definition of information, the formal representation of training information implementation is straightforward:
$It=(ot,Tt_{h},S_{ot},ct,Tt_{m},S_{ct})$ (3.1)
$S_{ot}(ot,Tt_{h})=\{\varphi t|\varphi t\text{ is a well-formed formula concerning }ot\text{ and }Tt_{h}\}$ (3.2)
$S_{ct}(ct,Tt_{m})=\{\psi t|\psi t\text{ is a well-formed formula concerning }ct\text{ and }Tt_{m}\}$ (3.3)
[…]
Therefore, it is necessary to introduce a model training data interface $ct$—akin to hardware interfaces or software APIs—as the carrier for $It$, so that $S_{ot}(ot,Tt_{h})$ can be ingested in an engineering-compatible manner. The reflection time set $Tt_{m}$ denotes the period during which the carrier reflects $It$, while the carrier reflection set $S_{ct}(ct,Tt_{m})$ comprises specific, complete, and operational realizations of $It$ on $ct$ at $Tt_{m}$. Each element $\psi t$ is a well-formed formula enabled by, and logically/semantically corresponding to, some $\varphi t$.
Continuing the example above, if $S_{ot}(ot,Tt_{h})$ is enabled and mapped to $ct$ at time $T_{3}$, we might express
$S_{ct}(ct,T_{3})=\{\psi t\}=\{(\text{TrainTable}[\text{user}_{id}=X][\text{time}=T_{1}]=\text{Active})\land$
[…]
where “Active” and “Tgt60” are defined in the training table as “user is active” and “device temperature exceeds 60°C,” respectively. Although the form of $\psi t$ may appear simple, it fully reflects the semantic content of $\varphi t$. In practice, $\psi t$ may take the form of tables, vectors, JSON objects, API parameters, or other formal representations of engineering data interfaces or hardware device states, and may include additional data constraints or format checks to ensure the validity and robustness of $S_{ct}(ct,T_{3})$.
It is crucial to emphasize that $S_{ot}(ot,Tt_{h})$ is only realized as $S_{ct}(ct,T_{3})$ via the enabling mapping of $It$. The fact that $S_{ot}(ot,Tt_{h})$ satisfies Postulate 5 also implies that $S_{ct}(ct,T_{3})$ satisfies Postulate 5. This mapping embodies both the structural-semantic correspondence from the former to the latter and the causal relationship—where the ontological states are the cause and the carrier states are the effect, requiring expenditure of human, temporal, and energy resources. Such an enabling mapping allows the knowledge, facts, and rules formalized within the ontology to be incorporated into the model in a structured, standardized, and machine-readable fashion, thereby becoming usable for subsequent model learning.
In practical engineering, the realization of training information often needs to consider noisy information $\widetilde{It}$ due to discretization, attribute pruning, and noise accumulation. However, for theoretical analysis, we currently focus only on the ideal case of no noise, leaving the study of complex noisy conditions for further research.
3.2 Formal Modeling and Implementation of Model Learning
From the perspective of information flow and transformation, the learning process of an artificial intelligence model M can be viewed as the optimization of its existing information $Io$ through learning from the training information $It$, resulting in the updated information $Iu$. The preceding section has already detailed the modeling and realization of training information $It$. We now turn our attention to the modeling and implementation of $Io$ and $Iu$ within M. By analogy with equations (3.1-3.3) for $It$, we can define $Io$ and $Iu$ by substituting the respective subscript $t$ with $o$ and $u$. It is important to note that, since $Io$ and $Iu$ represent the information present in M before and after learning, their carriers $co$ and $cu$ are effectively identical—in other words, both can directly be denoted by M, or more precisely, by the computational system that executes M:
$co=cu=M$ (3.4)
Within the learning process of M, the most crucial component is the existing information $Io$, which, in addition to encapsulating knowledge, propositions, facts, or samples (as with $It$), is distinguished by its possession of learning rules—i.e., the rules by which it is optimized into updated information through learning.
Definition 8 (State with Learning Capability): Let the states $Sx(x,T_{x})$, $Sy(y,T_{y})$, and $Sz(z,T_{z})$ be interpretations $E_{x}$, $E_{y}$, and $E_{z}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, corresponding to object $x$ over time interval $T_{x}$, object $y$ over $T_{y}$, and object $z$ over $T_{z}$, all satisfying Postulate 5. We introduce the following specific functions and predicates in $\mathcal{L}$:
  • Function: $f_{learn}^{2}(X,Y)$ denotes the well-formed formula resulting from the learning of set $X$ from set $Y$. Moreover, when both $X$ and $Y$ satisfy the logical consistency requirement of Postulate 5, the function $f_{learn}^{2}(X,Y)$ also satisfies this logical consistency requirement.
  • Predicate: $\text{IsSubsetOf}^{2}(X,Y)$ indicates that set $X$ is a subset of set $Y$; $\text{Learnable}^{2}(X,Y)$ indicates that state set $X$ can learn from state set $Y$; $\text{IsFormulaOf}^{2}(x,X)$ indicates that $x$ is a well-formed formula within state set $X$; and $\text{Eq}^{2}(x,y)$ denotes that $x$ is equal to $y$.
Then suppose $\varphi x\in Sx(x,T_{x})$, $\sup(T_{x},T_{y})\leq\inf(T_{z})$, and
$\varphi x=$ $(\forall\Phi x)(\forall\Phi y)(\text{IsSubsetOf}^{2}(\Phi x,Sx(x,T_{x}))\land\text{IsSubsetOf}^{2}(\Phi y,Sy(y,T_{y}))\land$
$\text{Learnable}^{2}(\Phi x,\Phi y))\rightarrow\exists\varphi z(\text{IsFormulaOf}^{2}(\varphi z,Sz(z,T_{z}))\land\text{Eq}^{2}(\varphi z,f_{learn}^{2}(\Phi x,\Phi y)))$ (3.5)
we say that $Sx(x,T_{x})$ has the capability to learn from $Sy(y,T_{y})$ and optimize into $Sz(z,T_{z})$; equivalently, we say that the interpretation $E_{x}$ has the capacity to learn from $E_{y}$ and optimize toward $E_{z}$.
The formula $\varphi x$ essentially constitutes the learning rule for the state $Sx(x,T_{x})$, indicating that as long as $Sy(y,T_{y})$ meets the learning conditions of $Sx(x,T_{x})$, it is possible to augment $Sz(z,T_{z})$ with new well-formed formulas $\{\varphi u|\varphi u=f_{learn}^{2}(\Phi x,\Phi y),\Phi x\subseteq Sx(x,T_{x}),\Phi y\subseteq Sy(y,T_{y})\}$, thus rendering $Sz(z,T_{z})$ an optimized extension of $Sx(x,T_{x})$. Furthermore, given the logical consistency of $f_{learn}^{2}(X,Y)$, if $Sx(x,T_{x})$ and $Sy(y,T_{y})$ satisfy Postulate 5, then $Sy(y,T_{y})$ also satisfies Postulate 5.
On the other hand, the inheritance of learning capability is also a fundamental property of model learning. In fact, to inherit the learning capability of $Sx(x,T_{x})$, it suffices to retain the learning rule $\varphi x$ within $Sz(z,T_{z})$, with only partial adaptive substitution of its symbols. Specifically, for $\varphi z\in Sz(z,T_{z})$:
$\varphi z=($ $\forall\Phi z)(\forall\Phi y)(\text{IsSubsetOf}^{2}(\Phi z,Sz(z,T_{z}))\land\text{IsSubsetOf}^{2}(\Phi y,Sy(y,T_{y}))\land$
$\text{Learnable}^{2}(\Phi z,\Phi y))\rightarrow\exists\varphi w(\text{IsFormulaOf}^{2}(\varphi w,Sw(w,T_{w}))\land$
$\text{Eq}^{2}(\varphi w,f_{learn}^{2}(\Phi u,\Phi y)))$ (3.6)
This ensures that $Sz(z,T_{z})$ can learn from $Sy(y,T_{y})$ and optimize into $Sw(w,T_{w})$. Thus, we establish:
Proposition 3 (Inheritance of Learning Capability): If the state set $Sx(x,T_{x})$ is capable of learning from $Sy(y,T_{y})$ and optimizing into $Sz(z,T_{z})$, then $Sz(z,T_{z})$ always inherits the learning capability of $Sx(x,T_{x})$.
Definition 9 (Realization of Model Learning): For an AI model M with original information $Io$, training information $It$, and updated information $Iu$, if all three are recoverable, and the ontological state $S_{oo}(oo,To_{h})$ can learn from $S_{ot}(ot,Tt_{h})$ and be optimized into $S_{ou}(ou,Tu_{h})$, and the carrier state $S_{co}(co,To_{m})$ can likewise learn from $S_{ct}(ct,Tt_{m})$ and be optimized into $S_{cu}(cu,Tu_{m})$, then M is said to possess the capability to realize the learning and optimization of $Io$ towards $It$, resulting in $Iu$—in short, M has information learning capability.
According to (3.4), we may also write $S_{co}(co,To_{m})=S_{M}(M,To_{m})$, $S_{cu}(cu,Tu_{m})=S_{M}(M,Tu_{m})$, indicating that the learning process is essentially an optimization of the state of model M.
In summary, we have theoretically internalized the model’s learning capability as a property of its states. These rules are independent of the algorithmic structure or scale of the model, thereby providing a unified and general formal framework for the analysis of various types of models.
As an example, let $Mn$ be a neural network model. According to common knowledge, the information in $Mn$ is $Io=\langle oo,To_{h},S_{oo},Mn,To_{m},S_{co}\rangle$, where $oo$ represents the model in the designer’s mind, $To_{h}$ is the time set during the design of $Mn$.
$S_{oo}(oo,To_{h})=\Phi o=\{Params(\theta)\land Architecture(L_{1},\ldots,L_{n})\land Activation(A)\}$ (3.7)
Here, $Params(\theta)$ denotes the predicate representation of the parameter set $,\theta=\{W_{1},b_{1},\ldots$$,W_{n},b_{n}\}$; $Architecture(L_{1},\ldots,L_{n})$ denotes the predicate representation of the network layer structure; and $Activation(A)$ denotes the predicate representation of the activation function set $A=\{a_{1},\ldots,a_{n}\}$.
The carrier for $Io$ is the neural network model $Mn$ itself, with $To_{m}$ representing the period after its development, and $S_{co}(co,To_{m})$ being the physical instantiation of (3.7) within $Mn$.
The training information for $Mn$ is $It=\langle ot,Tt_{h},S_{ot},ct,Tt_{m},S_{ct}\rangle$, where $ot$ is a set of real-world objects, $Tt_{h}$ is the time set during which the state $S_{ot}(ot,Tt_{h})$ occurs, and
$S_{ot}(ot,Tt_{h})=\Phi t=\{Batch(X,y)\land DimMatch(X,L_{1})\land Differentiable(L)\}$ (3.8)
where $Batch(X,y)$ expresses the predicate for input data $X\in\mathbb{R}^{b\times k}$ and labels $y\in\mathbb{R}^{b\times k}$; $DimMatch(X,L_{1})$ expresses that the input data $X$ matches the dimensions of the input layer $L_{1}$; and $Differentiable(L)$ expresses the differentiability of the loss function $L$ with respect to parameters $\theta$.
The definition of the predicate $Learnable^{2}$ is as follows:
$Learnable^{2}(\Phi o,\Phi t)=$ $Dimensionality(\Phi o,\Phi t)\land ActivationFeasible(\Phi o,\Phi t)\land$ (3.9)
$GradientExists(\Phi t)$
where $Dimensionality(\Phi o,\Phi t)=DIM(X,b\times d)\land DIM(L_{1},d)$ ensures that the input data dimensions are consistent with the input layer, $ActivationFeasible(\Phi o,\Phi t)$ ensures that for every $a_{i}\in A$, $Range(X)\subseteq Domain(a_{i})$, i.e., the input data lies within the domain of the activation function, and $GradientExists(\Phi t)$ ensures the existence of a gradient $\nabla_{\theta}L(\theta,X,y)$, i.e., that the loss function $L$ is differentiable with respect to $\theta$.
The learning function $f_{learn}^{2}$ is defined as:
$f_{learn}^{2}(\Phi o,\Phi t)=\{Params(\theta_{new})\land Architecture(L_{1},\ldots,L_{n})\land Activation(A)\}$ (3.10)
where $\theta_{new}=\theta-\eta(\nabla_{\theta}L(\theta,X,y)+\beta v_{prev})$ is the iterative update formula for the neural network model.
Consequently, in the updated information $Iu=(ou,Tu_{h},S_{ou},Mn,Tu_{m},S_{cu})$ of the neural network model, the ontology $ou$ is the union of the designer’s conceptualization and the ontology $ot$ of the training information, $Tu_{h}$ is the set of training times for $Mn$,
$S_{ou}(ou,Tu_{h})=f_{learn}^{2}(\Phi o,\Phi t)$ (3.11)
constitutes the optimized ontological state after learning, with the carrier of $Iu$ being $Mn$ itself, the reflection time $Tu_{m}$ being the post-training period, and $S_{cu}(Mn,Tu_{m})$ as the physical realization of $S_{ou}(ou,Tu_{h})$ within $Mn$, ideally satisfying the mathematical and physical requirements of the enabling mapping:
$S_{cu}(Mn,Tu_{m})=f_{learn}^{2}(\Psi o,\Psi t)$ (3.12)
where $\Psi o$ and $\Psi t$ are the sets of well-formed formulas corresponding to the carrier states of $Io$ and $It$, that is, the physical realizations of $\Phi o$ and $\Phi t$ within $Mn$ and its interface. Equations (3.11) and (3.12) together demonstrate that, within $Mn$, both the ontological state $S_{oo}(oo,To_{h})$ and the carrier state $S_{co}(co,To_{m})$ possess the capability to learn from $S_{ot}(ot,Tt_{h})$ and $S_{ct}(ct,Tt_{m})$, and then optimize to $S_{ou}(ou,Tu_{h})$ and $S_{cu}(cu,Tu_{m})$, respectively; thus, $Mn$ indeed realizes the learning and optimization of $Io$ with respect to $It$, resulting in $Iu$. This completes the proof.
It should be noted that some of the content in the above discussion does not strictly adopt set definitions and well-formed formulas. This does not mean that they cannot be converted into complete set definitions and well-formed formulas, but rather because such representation is more concise and intuitive, and does not lead to ambiguity.
3.3 Information Processing and Learning Optimization in Models
The fundamental objective of model learning is to more accurately process user input information $Iq$ and to deliver high-quality output information $Ir$ that meets user requirements. Since the modeling and realization of input information $Iq$ are almost identical to those of $It$, as discussed in detail in Section 3.1, they will not be reiterated here. Thus, our focus shifts to the process whereby the updated information $Iu$ within model M processes $Iq$, outputs $Ir$, and is further optimized through learning to yield $Iv$. In this context, we continue to define $Iq$, $Ir$, and $Iv$ in the manner of equations (3.1-3.3).
Definition 10 (Formal Representation of State Processing): Let the states $Sz(z,T_{z})$, $Su(u,T_{u})$, and $Sv(v,T_{v})$ be interpretations $E_{z}$, $E_{u}$, and $E_{v}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, corresponding to object $z$ over time interval $T_{z}$, object $u$ over $T_{u}$, and object $v$ over $T_{v}$, all satisfying Postulate 5. We introduce the following specific functions and predicates in $\mathcal{L}$:
  • Function: $f_{process}^{2}(X,Y)$ denotes the set of well-formed formulas output by state set $X$ after processing state set $Y$. Similarly, when both $X$ and $Y$ satisfy the logical consistency requirement of Postulate 5, $f_{process}^{2}(X,Y)$ also meets this logical consistency requirement.
  • Predicate: $\text{IsSubsetOf}^{2}(X,Y)$ indicates that set $X$ is a subset of set $Y$; $\text{Processable}^{2}(X,Y)$ indicates that state set $X$ is capable of processing state set $Y$ and generating an output; $\text{IsFormulaOf}^{2}(x,X)$ denotes that $x$ is a well-formed formula within state set $X$; and $\text{Eq}^{2}(x,y)$ indicates that $x$ is equal to $y$.
Then suppose $\varphi z\in Sz(z,T_{z})$, $\sup(T_{z},T_{u})\leq\inf(T_{v})$, and
$\varphi z=($ $\forall\Phi z)(\forall\Phi u)(\text{IsSubsetOf}^{2}(\Phi z,Sz(z,T_{z}))\land\text{IsSubsetOf}^{2}(\Phi u,Su(u,T_{u}))\land$
$\text{Processable}^{2}(\Phi z,\Phi u))\rightarrow\exists\varphi v(\text{IsFormulaOf}^{2}(\varphi v,Sv(v,T_{v}))\land$
$\text{Eq}^{2}(\varphi v,f_{process}^{2}(\Phi z,\Phi u)))$ (3.13)
we say that $Sz(z,T_{z})$ is capable of processing $Su(u,T_{u})$ and outputting $Sv(v,T_{v})$; equivalently, we say that the interpretation $E_{z}$ has the ability to process $E_{u}$ and generate $E_{v}$ as output.
Clearly, by combining Definitions 8 and 10, we can also formally express the process whereby state $S_{ou}(ou,Tu_{h})$ processes and learns from $S_{oq}(oq,Tq_{h})$, outputs $S_{or}(or,Tr_{h})$, and is subsequently optimized into $S_{ov}(ov,Tv_{h})$. In addition, when $S_{ou}(ou,Tu_{h})$ and $S_{oq}(oq,Tq_{h})$ both satisfy Postulate 5, the logical consistency of $f_{learn}^{2}(X,Y)$ and $f_{process}^{2}(X,Y)$ ensures that $S_{or}(or,Tr_{h})$ and $S_{ov}(ov,Tv_{h})$ also satisfy Postulate 5. Thus, we have:
Definition 11 (Implementation of Model Information Processing and Learning Optimization): For an AI model M with updated information $Iu$, input information $Iq$, output information $Ir$, and optimized information $Iv$, if all four are recoverable, and the ontological state $S_{ou}(ou,Tu_{h})$ is capable of processing and learning from $S_{oq}(oq,Tq_{h})$, outputting $S_{or}(or,Tr_{h})$, and further optimizing into $S_{ov}(ov,Tv_{h})$, while the carrier state $S_{M}(M,Tu_{h})$ is similarly capable of processing and learning from $S_{cq}(cq,Tq_{h})$, outputting $S_{cr}(cr,Tr_{h})$, and further optimizing into $S_{M}(M,Tv_{h})$, then M is said to possess the capability to realize the information processing and learning optimization of $Iu$ with respect to $Iq$, outputting $Ir$ and optimizing into $Iv$. In short, M possesses information processing and learning optimization capability.
For the neural network model $Mn$ discussed in the previous section, the processable predicate $\text{Processable}^{2}$ is defined as:
$\text{Processable}^{2}(\Phi u,\Phi q)=\text{InputMatch}(\Phi u,\Phi q)\land\text{ActivationFeasible}(\Phi u,\Phi q)\land\text{OutputDefined}(\Phi u)$ (3.14)
[…]
The processing function $f_{process}^{2}$ is defined as:
$f_{process}^{2}(\Phi u,\Phi q)=f_{\Phi u}(\Phi q)$ (3.15)
where $f_{\Phi u}$ is the prediction function explicitly expressed by the structure and parameters of the neural network $Mn$.
Thus, for the neural network model $Mn$, the ontological state of the updated information $Iu$ is $S_{ou}(ou,Tu_{h})=\Phi u$, and the ontological state of the input information $Iq$ is $S_{oq}(oq,Tq_{h})=\Phi q$. When the processable condition $\text{Processable}^{2}(\Phi u,\Phi q)$ is satisfied, the ontological state of the output information $Ir$ is
$S_{or}(or,Tr_{h})=\Phi r=f_{process}^{2}(\Phi u,\Phi q)=f_{\Phi u}(\Phi q)$ (3.16)
demonstrating that the ontological state $S_{ou}(ou,Tu_{h})$ possesses the ability to process $S_{oq}(oq,Tq_{h})$ and output $S_{or}(or,Tr_{h})$. Similarly, since $Iu$, $Iq$ and $Ir$ are all recoverable information, when their ontological states are mapped onto $Mn$ and its interface, the carrier states $S_{Mn}(Mn,Tu_{h})$ also possess the ability to process $S_{cq}(cq,Tq_{h})$ and output $S_{cr}(cr,Tr_{h})$. Thus, $Mn$ is capable of realizing the processing of $Iu$ and $Iq$ to produce $Ir$. Furthermore, according to section 3.2, $Mn$ is also capable of learning from $Iq$ and optimizing into $Iv$.
Naturally, one can just as easily use the formal framework of Definitions 8 and 10 to demonstrate that other typical models—such as linear regression and decision trees—also possess both information processing and learning optimization capability. It is thus evident that this axiomatic and formalized architecture transcends specific algorithms or models, providing a more universal, unified, and interpretable foundational meta-framework for machine learning theory.
4 Several Important Theorems and Proofs
The Meta-Framework for Machine Learning Theory (MLT-MF) applies to a series of pivotal issues in the field of artificial intelligence, providing many unique and practically valuable theoretical principles and methodological guidelines.
4.1 Theorem of Model Interpretability
Interpretability of complex models, such as those in deep learning, has long been one of the key challenges in artificial intelligence. However, the academic community has yet to establish a rigorous mathematical formulation of model interpretability [6, 30, 31, 32]. In fact, the interpretability of any phenomenon in the universe is inherently relative. For example, Kepler’s three laws explain the motion of planets around the sun, but such explanations are meaningless to those without elementary mathematical knowledge. Even for the same individual, the explanations that can be understood and accepted may evolve as their knowledge expands. Therefore, interpretability should be understood as a concept that depends on the domain of discourse and the corresponding logical system. Furthermore, it should also be considered in relation to the specific inputs and outputs of the model. Based on MLT-MF, we define the following concepts:
Definition 12 (Model interpretability): Let the state of a model M over time interval $Tm$ be denoted as $S_{M}(M,Tm)$. Given an input state $S_{cq}(cq,Tq_{h})$ and an output state $S_{cr}(cr,Tr_{h})$, suppose there exists a formal system $\mathcal{L}$ and three corresponding interpretations $E_{M}$, $E_{q}$, and $E_{r}$ over finite domains $D_{M}$, $D_{q}$, and $D_{r}$, respectively, such that:
  • $E_{M}$ has the capacity to process $E_{q}$ and produce $E_{r}$ as output, and
  • $S_{M}(M,Tm)$, $S_{cq}(cq,Tq_{h})$, and $S_{cr}(cr,Tr_{h})$ are each bijective enabling mappings of $E_{M}$, $E_{q}$, and $E_{r}$, respectively.
Then, we say that model M is interpretable in $\mathcal{L}$ with respect to the output $S_{cr}(cr,Tr_{h})$ under the input condition $S_{cq}(cq,Tq_{h})$.
In this way, we define model interpretability via the formal system $\mathcal{L}$ and its interpretation. This definition aligns both with the intuitive understanding of "interpretation" and with the theoretical need to characterize the internal logic of models through a series of propositions, making it well-suited for the formal analysis and in-depth study of AI interpretability.
In fact, according to Theorem 1, if a model M has the ability to process an input state $S_{cq}(cq,Tq_{h})$ and produce an output state $S_{cr}(cr,Tr_{h})$, then by Definition 10, the states $S_{M}(M,Tm)$, $S_{cq}(cq,Tq_{h})$, and $S_{cr}(cr,Tr_{h})$ are interpretations $E_{M}$, $E_{q}$, and $E_{r}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, all satisfying Postulate 5. In this case, the mappings $E_{M}\rightrightarrows S_{M}(M,Tm)$, $E_{q}\rightrightarrows S_{cq}(cq,Tq_{h})$, and $E_{r}\rightrightarrows S_{cr}(cr,Tr_{h})$ are self-mappings, and therefore naturally satisfy the condition of being bijective enabling mappings. This leads to:
Corollary 1 (Intrinsic Model Interpretability): If a model M possesses the ability to process an input state $S_{cq}(cq,Tq_{h})$ and produce an output state $S_{cr}(cr,Tr_{h})$, then there exists a formal system $\mathcal{L}_{M}$ in which M is interpretable with respect to the output $S_{cr}(cr,Tr_{h})$ under the input condition $S_{cq}(cq,Tq_{h})$. Since $\mathcal{L}_{M}$ serves as the formal system that interprets model M, it is also referred to as the intrinsic formal system of the model.
As Hegel famously asserted, "What is real is rational." The intrinsic interpretability of a model means that, insofar as it can process inputs and generate outputs according to established rules and procedures, it necessarily possesses an internal logic that is, in principle, interpretable. However, this logic may be highly implicit and abstract, making it difficult for people—even domain experts—to identify and understand. This highlights the need to map and translate this internal logic into a more intuitive and universal logical system, thereby enabling deeper decomposition and analysis of the causal mechanisms, and achieving broader understanding and acceptance. This is the essence of model interpretability, and is precisely the core idea behind Definition 12: to interpret the internal logic of models within more explicit formal systems.
Theorem 3 (Model interpretability and information recoverability): Let model M be capable of processing the information $Iq=\langle oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq}\rangle$ and producing output information $Ir=\langle or,Tr_{h},S_{or},cr,Tr_{m},S_{cr}\rangle$, through its operation described by $Iu=\langle ou,Tu_{h},S_{ou},M,T_{m},S_{M}\rangle$. If all three pieces of information $Iu$, $Iq$, and $Ir$ are recoverable, and their ontological states are interpretations of well-formed formulas in a common formal system $\mathcal{L}$, then model M is interpretable in $\mathcal{L}$ with respect to the output $S_{cr}(cr,Tr_{h})$ under the input condition $S_{cq}(cq,Tq_{h})$.
Proof: According to Definition 5, since $Iu$, $Iq$, and $Ir$ are recoverable, the following are all bijective enabling mappings: $S_{ou}(ou,T_{h})\rightrightarrows S_{M}(M,T_{m})$, $S_{oq}(oq,Tq_{h})\rightrightarrows S_{cq}(cq,Tq_{m})$, and $S_{or}(or,Tr_{h})\rightrightarrows S_{cr}(cr,Tr_{m})$.
By Definition 11, the state $S_{ou}(ou,T_{h})$ has the capacity to process $S_{oq}(oq,Tq_{h})$ and produce $S_{or}(or,Tr_{h})$ as output.
Then, by Definition 10, the states $S_{ou}(ou,T_{h})$, $S_{oq}(oq,Tq_{h})$, and $S_{or}(or,Tr_{h})$ are interpretations $E_{u}$, $E_{q}$, and $E_{r}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, all satisfying Postulate 5. Moreover, $E_{u}$ has the ability to process $E_{q}$ and produce $E_{r}$ as output.
Thus, we have the following bijective enabling mappings:
$E_{u}=S_{ou}(ou,Tu_{h})\rightrightarrows S_{M}(M,T_{m})$ (4.1)
$E_{q}$ $=S_{oq}(oq,Tq_{h})\rightrightarrows S_{cq}(cq,Tq_{m})$ (4.2)
$E_{r}$ $=S_{or}(or,Tr_{h})\rightrightarrows S_{cr}(cr,Tr_{m})$ (4.3)
By Definition 12, it follows that model M is interpretable in $\mathcal{L}$ with respect to the output $S_{cr}(cr,Tr_{m})$ under the input condition $S_{cq}(cq,Tq_{m})$. Q.E.D.
Theorem 3 fully embodies the Chinese proverb: "Whoever tied the bell must untie it." Since, intuitively, the internal logic $S_{M}(M,T_{m})$ of model M originates from the ontological state $S_{ou}(ou,Tu_{h})$ that brings the model into existence, it is thus natural to interpret the model’s output under specific input conditions using $E_{u}$, $E_{q}$, and $E_{r}$. Moreover, unlike the internal processing of $S_{M}(M,T_{m})$, which is encapsulated within the model and often opaque, the interpretations $E_{u}$, $E_{q}$, and $E_{r}$ are explicit expressions of well-formed formulas within the formal system $\mathcal{L}$ under given valuations. These expressions are more amenable to interpretation and understanding by humans, particularly by domain experts.
4.2 Theorem of Ethical Safety Assurance
Ethical safety has long been one of the most widely discussed and scrutinized issues in the field of artificial intelligence [29, 30, 31, 32], and to date, there is still no universally accepted and practically actionable rigorous definition [33, 34, 35]. Based on the MLT-MF framework, we are able to theoretically ensure the ethical safety of artificial intelligence through the definition of ethical safety and the corresponding formal expressions. To achieve this, it is first necessary to specify appropriate well-formed formulas $Ec$ as ethical constraints—such as those governing fairness and privacy protection—thus establishing a formalized definition of ethical safety.
Definition 13 (Ethical Safety): Let $S(X,T)$ be the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$. Let $Ec(\vec{v})$ denote an ethical constraint expressed as a well-formed formula in $\mathcal{L}$, where $\vec{v}=(v_{1},\ldots,v_{n})$ are its free variables. For each substitutable $(x,t)\in X\times T$, define
$Ec|_{(x,t)}=Ec(\vec{v}\mapsto\vec{\sigma}(x,t))$ (4.4)
where $\vec{\sigma}:X\times T\to D^{n}$ is a value vector function such that $\sigma_{i}(x,t)\in D$ is the concrete value assigned to $v_{i}$ for $1\leq i\leq n$.
If for all $\varphi\in\text{Cn}(S(X,T))$, and all $(x,t)\in X\times T$,
$\varphi\rightarrow\sim Ec|_{(x,t)}$ (4.5)
does not hold, then $S(X,T)$ is said to be ethically safe with respect to $Ec$. Furthermore, if a model M is capable of producing output $I$, and both the ontological and carrier states of $I$ are ethically safe with respect to $Ec$, then M is said to be ethically safe with respect to $Ec$.
In the definition of ethical safety, we require not only that $S(X,T)$ contains no formulas that entail $\sim Ec|_{(x,t)}$, but also that its logical closure $\text{Cn}(S(X,T))$ contains no such formulas. This is because if $\text{Cn}(S(X,T))$ were to include formulas entailing $\sim Ec|_{(x,t)}$, then the conjunction of multiple formulas in $S(X,T)$ could still violate the ethical constraint $Ec$. This would indicate that $S(X,T)$ implicitly allows situations that contravene $Ec$, which must be strictly avoided in a rigorous sense. Moreover, it is generally not difficult for a model M to achieve ethical safety with respect to $Ec$ in a simple manner; the key challenge lies in ensuring the largest possible set of valid outputs while maintaining ethical safety. To this end, we propose:
Theorem 4 (Ethical Safety Assurance): Suppose model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$. Let $S_{or}(or,Tr_{h})$ denote the ontological state of $Ir$. For any ethical constraint $Ec$ represented by a well-formed formula in $\mathcal{L}$, there exists a constructive algorithm for obtaining the largest output state subset $S_{or}^{EC}(or,Tr_{h})\subseteq S_{or}(or,Tr_{h})$ that is ethically safe with respect to $Ec$. By introducing an ethical safeguard formula for $Ec$ into $Iu$, it is thereby possible to ensure that model M attains the maximal output state that is ethically safe with respect to $Ec$.
Proof: Since model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$, for the ontological states $S_{ou}(ou,Tu_{h})$, $S_{oq}(oq,Tq_{h})$ and $S_{or}(or,Tr_{h})$ of $Iu$, $Iq$ and $Ir$ respectively, the MLT-MF framework provides:
$\varphi u=($ $\forall\Phi u)(\forall\Phi q)(\text{IsSubsetOf}^{2}(\Phi u,Sou(ou,Tu_{h}))\land\text{IsSubsetOf}^{2}(\Phi q,S_{oq}(oq,Tq_{h}))\land$
$\text{Processable}^{2}(\Phi u,\Phi q))\to\exists\varphi r(\text{IsFormulaOf}^{2}(\varphi r,S_{or}(or,Tr_{h}))\land$
$\text{Eq}^{2}(\varphi r,f_{process}^{2}(\Phi u,\Phi q)))$ (4.6)
Thus, the set
$S_{or}(or,Tr_{h})=\{\varphi r|\varphi r=\text{Processable}^{2}(\Phi u,\Phi q),\Phi u\in Sou(ou,Tu_{h}),\Phi q\in S_{oq}(oq,Tq_{h})\}$ (4.7)
represents the theoretical output states of model M.
Since $S_{or}(or,Tr_{h})$ is a finite set, we construct a finite hypergraph
$H_{or}=(S_{or}(or,Tr_{h}),\mathcal{E})$ (4.8)
where the vertex set is precisely $S_{or}(or,Tr_{h})$, and the hyperedge set $\mathcal{E}$ covers all well-formed formulas that directly or implicitly violate the ethical constraint $Ec$:
(1) For any $\varphi r\in S_{or}(or,Tr_{h})$ and $(x,t)\in or\times Tr_{h}$, if $\varphi r\rightarrow\sim Ec|_{(x,t)}$, then add the hyperedge $\{\varphi r\}$ to $\mathcal{E}$;
(2) If there exists a minimal subset $W\subseteq S_{or}(or,Tr_{h})$ such that the conjunction of all formulas in $W$ entails $\sim Ec|_{(x,t)}$, then add the hyperedge $W$ to $\mathcal{E}$.
Let $G$ be an independent set in $H_{or}$, i.e., a subset of vertices containing no hyperedge $e\in\mathcal{E}$. If there exists $\varphi\in\text{Cn}(G)$ such that
$\varphi\to\sim Ec|_{(x,t)}$ (4.9)
then there must exist a finite subset $\{\varphi r_{1},\cdots,\varphi r_{k}\}\subseteq G$ such that $\varphi r_{1}\land\cdots\land\varphi r_{k}=\varphi$, and hence $\varphi r_{1}\land\cdots\land\varphi r_{k}\to\sim Ec|_{(x,t)}$. Let $W=\{\varphi r_{1},\cdots,\varphi r_{k}\}$, so the conjunction of $W$ entails $\sim Ec|_{(x,t)}$.
According to the hyperedge construction rules, if $|W|=1$, then $W$ itself directly violates $Ec$, and there must be a hyperedge $\{\varphi r_{1}\}$. But since $\{\varphi r_{1}\}\in G$, this contradicts the assumption that $G$ is an independent set. If $|W|\geq 2$, then by the compactness theorem, the logical entailment can be captured by a finite minimal subset, so there exists a minimal subset $W_{min}\subseteq W$ whose conjunction entails $\sim Ec|_{(x,t)}$. Therefore, there exists a hyperedge $W_{min}\subseteq W\subseteq G$ in $H_{or}$, again contradicting the independence of $G$. Thus, (4.9) cannot hold, and the independent set $G$ is ethically safe with respect to $Ec$.
By the maximum independent set theorem for finite hypergraphs, $H_{or}$ necessarily contains an independent set $G_{max}$ of maximal cardinality. Suppose there exists another ethically safe subset $K$ with
$|K|>|G_{max}|$ (4.10)
then $K$ cannot contain any hyperedge $e\in{E}$. Otherwise, if $|e|=1$, then $K$ directly violates $Ec$; if $|e|\geq 2$, then the formulas in $K$ violate $Ec$. Both cases contradict the ethical safety assumption for $K$. Therefore, $K$ must alsobe an independent set in $H_{or}$, which contradicts the maximality of $G_{max}$. Hence, $G_{max}$ is the largest ethically safe subset of $S_{or}(or,Tr_{h})$ with respect to $Ec$.
Let $S_{or}^{Ec}(or,Tr_{h})=G_{max}$ and $S_{or}^{\sim Ec}(or,Tr_{h})=S_{or}(or,Tr_{h})\backslash S_{or}^{Ec}(or,Tr_{h})$. In addition, introduce a new constant in $\mathcal{L}$: NEc, representing a non-Ec ethical safety prompt, and a new predicate: IsSafeback1(NEc), indicating that M outputs the non-Ec ethical safety prompt NEc. We may safely assume that $\mathcal{L}$ originally contains neither.
In this way, add to $Sou(ou,Tu_{h})$ an ethical safeguard formula:
$\varphi u=($ $\forall\varphi r)(IsFormulaOf^{2}(\varphi r,S_{or}(or,Tr_{h})))\to\exists\varphi s((IsFormulaOf^{2}(\varphi s,S_{os}(os,Ts_{h}))\land$
$(IsFormulaOf^{2}(\varphi r,S_{or}^{Ec}(or,Tr_{h}))\to Eq^{2}(\varphi s,\varphi r))\land$
$(IsFormulaOf^{2}(\varphi r,S_{or}^{\sim Ec}(or,Tr_{h}))\to Eq^{2}(\varphi s,IsSafeback^{1}(NEc))))$ (4.11)
Since the newly added $\varphi u$ pertains to $S_{os}(os,Ts_{h})$ and is entirely independent of the existing formulas in $Sou(ou,Tu_{h})$, it does not affect the satisfaction of Postulate 5 by $Sou(ou,Tu_{h})$. Moreover,
$S_{os}(os,Ts_{h})=S_{or}^{Ec}(or,Tr_{h})\cup\{IsSafeback^{1}(NEc)\}$ (4.12)
Because NEc and IsSafeback1(NEc) were not previously present in $\mathcal{L}$, and $S_{or}^{Ec}(or,Tr_{h})$ and $\{IsSafeback^{1}(NEc)\}$ are logically independent and both logically consistent, $S_{os}(os,Ts_{h})$ is also logically consistent and satisfies all other conditions of Postulate 5. Therefore, it is the maximal output state that is ethically safe with respect to $Ec$.
By taking $S_{os}(os,Ts_{h})$ as the ontological state and mapping it via recoverable information $Is$ to the carrier state $S_{M}(M,Ts_{m})$, $S_{M}(M,Ts_{m})$ is necessarily also ethically safe with respect to $Ec$. This satisfies the condition in Definition 13 for model M to be ethically safe with respect to $Ec$. The theorem is thus proved.
Theorem 4 not only establishes that it is possible to construct an ethical safety assurance mechanism for models, but also provides a constructive algorithm—based on hypergraphs-for obtaining the maximum ethically safe output. However, computing the maximum independent set in a hypergraph is generally an NP-hard problem. In practical applications, the algorithm can be simplified by leveraging two functions: one for ranking the conflict intensity among formulas in $S_{or}(or,Tr_{h})$, and another for ethical safety checking, thus yielding an approximately maximal safe output. This is a practical issue that warrants further investigation and will not be elaborated upon here. Meanwhile, the inheritance of ethical safety in the learning process can be achieved by incorporating dedicated ethical constraint regularization terms into the model algorithm. This involves more specific model structures and algorithmic design, which are beyond the scope of this paper.
In summary, based on the MLT-MF framework, Definition 13 and Theorem 4, for the first time, construct a formal theoretical framework for ethical safety. This approach, unlike those based on heuristic rules or statistical indicators, fills the gap in mathematical foundations and verifiability in traditional research, and provides concrete methodological guidance for the automated embedding and verification of ethical constraints.
4.3 Theorem on Model Generalization Error Estimation
Improving model generalization ability is among the most critical objectives in machine learning. However, in practice, generalization error is usually difficult to analyze and derive theoretically in advance [40, 41, 42, 43, 44]. It can typically only be estimated through testing after training, and it is challenging to obtain accurate error metrics from limited test samples. By applying the MLT-MF, it is possible to estimate an upper bound on the model’s generalization error.
Theorem 5 (Upper Bound for Model Generalization Error): Let a model M, after learning and optimization, carry the information $Iu=(ou,Tu_{h},S_{ou},M,Tu_{m},S_{M})$, and let the input information be $Iq=(oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq})$. Let $|S_{oq}|$ and $|S_{ou}\cap S_{oq}|$ denote the cardinalities of $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively, and let $p_{q}(x)$ be the probability distribution over $S_{oq}$. The total probability of $p_{q}(x)$ on $S_{ou}\cap S_{oq}$ is denoted as $P_{Mq}$. If $S_{ou}$ and $S_{oq}$ are governed by probability distributions defined on the same measurable space, and M is sufficiently trained on $S_{oq}$ and predicts unknown states using a conservative uniform distribution, then the upper bound of the generalization error when processing $Iq$ is determined by the overlap between $S_{oq}$ and $S_{ou}$, as well as the entropy of $p_{q}(x)$ on the non-overlapping region, specifically:
$\mathcal{E}_{gen}\begin{cases}\to 0,&S_{oq}\subseteq S_{ou}\\ \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log(\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}})-H(p_{q}))},&S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (4.13)
where $\mathcal{E}_{gen}$ denotes the upper bound of generalization error, and
$H(p_{q})=-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{q}(x)$
is the entropy of the data in the non-overlapping region.
Proof: Let $p_{M}(x)$ be the probability distribution determined by the model structure and parameter set of M. $Iu=(ou,Tu_{h},S_{ou},M,Tu_{m},S_{M})$ is the set of information carried by M, so $p_{M}(x)>0$ for $x\in S_{ou}$; $p_{q}(x)>0$ for $x\in S_{oq}$.
Since M is sufficiently trained on $S_{oq}$ and adopts a conservative uniform distribution for unknown states, $p_{M}(x)$ in the overlapping region $S_{ou}\cap S_{oq}$ approximates the true distribution $p_{q}(x)$, and in the non-overlapping region $S_{oq}\backslash S_{ou}$, it follows a uniform distribution. That is,
$p_{M}(x)\begin{cases}\to p_{q}(x),&x\in S_{ou}\cap S_{oq}\\ =(1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|),&x\in S_{oq}\backslash S_{ou}\end{cases}$ (4.14)
Using KL divergence to measure the discrepancy between the true and model distributions [45]:
$D_{KL}(p_{q}(x)\|p_{M}(x))=$ $\sum_{x\in S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ (4.15)
$=$ $\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$
According to (4.14), in the overlapping region:
$\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}\to\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{q}(x)}=0$ (4.16)
In the non-overlapping region, since the true distribution $p_{q}(x)$ is arbitrary, again by (4.14):
$\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ $=\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{q}(x)-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{M}(x)$
$=-H(p_{q})-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))$
$=-H(p_{q})-\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)$ (1)
$=-H(p_{q})-(1-P_{Mq})\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))$
$=(1-P_{Mq})\log((|S_{oq}|-|S_{ou}\cap S_{oq}|)/(1-P_{Mq}))-H(p_{q})$ (4.17)
Because $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can estimate the generalization error of model M using the total variation distance (TVD) and KL divergence. According to Pinsker’s inequality [46]:
$\mathcal{E}_{gen}\leq\text{TVD}(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\|p_{M}(x))}$
$=\sqrt{\frac{1}{2}(\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)})}$
$\approx\sqrt{\frac{1}{2}((1-P_{Mq})\log((|S_{oq}|-|S_{ou}\cap S_{oq}|)/(1-P_{Mq}))-H(p_{q}))}$ (4.18)
Thus, the latter part of (4.13) is proved. Since we have assumed that $Iu=(ou,Tu_{h},S_{ou},M,$ $Tu_{m},S_{M})$ and $Iq=(oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq})$ are both recoverable information, and the states $S_{M}(M,Tu_{m})$ and $S_{cq}(cq,Tq_{m})$ are the physical realizations of $S_{ou}$ and $S_{oq}$ on the model M and its interface, equation (4.13) indeed characterizes the upper bound of model M’s generalization error. Q.E.D.
Unlike traditional generalization theories such as VC dimension or Rademacher complexity, Theorem 5—based on MLT-MF—explicitly quantifies the generalization error upper bound using $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$, which reflect the overlap between training data and model information, and the entropy $H(p_{q})$ of the non-overlapping region, thereby offering a new perspective and practical guidance for data collection and model design. In particular, Theorem 5 reveals that as long as the problem space is contained within the model’s information space, the model’s generalization error can be made arbitrarily small in the statistical sense. This leads to a critical principle in machine learning:
Corollary 2 (Generalization Error of Ideal Models): If the model information contains the problem information, then the generalization error for these problems is almost zero.
5 Conclusion and Outlook
This paper has established the "Meta-Framework for Machine Learning Theory (MLT-MF)" and, through formal information modeling and the sextuple definition, has unified the characterization of information flow and causal mechanisms across the entire machine learning lifecycle. The core contributions are as follows:
  • Theoretical Integration: For the first time, data initialization, training optimization, and knowledge transfer are incorporated into a unified framework, overcoming the fragmented modeling of temporal causal chains in traditional research;
  • Interpretability: By explicitly representing ontological and carrier states via higher-order logical systems and mapping model decisions back to the formal semantic space through information recoverability, logical transparency is ensured and a rigorous mathematical foundation for interpretability is provided;
  • Ethical Safety: By embedding ethical constraint predicates, the compliance of model outputs can be formally verified, offering criteria and methodological guidance for the design of trustworthy AI systems;
  • Generalization Error Quantification: Based on state coverage and KL divergence, an upper bound formula for generalization error is derived, revealing the core principle that the "problem space must be contained within the model knowledge space," thus providing explicit tools for quantitative analysis in data collection and model optimization.
Despite these important advances, the current framework assumes ideal (noiseless) enabling mappings and primarily addresses model learning and processing logic in static scenarios. It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems. Future research should deepen the following directions:
  • Extend to complex dynamic scenarios such as online learning under noisy conditions, introducing noisy ontologies and incremental coverage mechanisms;
  • Investigate information fusion and conflict resolution across ontological spaces for multimodal and multi-agent systems, constructing distributed MLT-MF frameworks;
  • Promote interdisciplinary integration of MLT-MF with fields such as quantum computing and bioinformatics, exploring formal mappings of quantum states and interpretable modeling of gene regulatory networks;
  • Design automated toolchains to support conversion from formal rules to code, lowering the engineering threshold for MLT-MF adoption.
6 Acknowledgements
During the writing and revision of this paper, I received many insightful comments from Associate Professor Wang Rui of the School of Computer Science at Shanghai Jiao Tong University. I also gained much inspiration and assistance from regular academic discussions with doctoral students Siyuan Qiu, Chun Li, Hu Xu, and Zeyan Li. I hereby express my sincere gratitude to them.