### v9

# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

Abstract
[Objective] This study addresses key challenges in machine learning, namely the absence of a unified formal theoretical framework and the lack of foundational theories for model interpretability and ethical safety.
[Methods] We first construct a formal information model, explicitly defining the ontological states and carrier mappings of typical machine learning stages using sets of well-formed formulas. By introducing learnable and processable predicates, as well as learning and processing functions, we analyze the causal chain logic and constraint laws governing machine learning processes.
[Results] We establish the Machine Learning Theory Meta-Framework (MLT-MF), on which we further propose universal definitions for model interpretability and ethical safety. We prove and validate three key theorems: the relationship between model interpretability and information existence, ethical safety assurance, and the upper bound estimation of total variation distance (TVD).
[Limitations] The current framework assumes ideal, noise-free information enabling mappings and focuses primarily on model learning and processing logic in static scenarios. It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems.
[Conclusions] This work overcomes the limitations of fragmented research and provides a unified theoretical foundation for systematically addressing critical issues in contemporary machine learning.
Keywords: machine learning; objective information theory; formal logic systems; interpretability; ethical and safety assurance; generalization error
1 Introduction
Although artificial intelligence is currently the most dynamic field within information science and even the broader domain of science and technology, existing research still faces several significant shortcomings:
First, there is a lack of a formalized framework that covers the entire lifecycle of machine learning Rob2021 ; Marius2023 ; Meng2023 ; Alquier2024 ; Vilalta2002 . Current theoretical studies in machine learning tend to focus on isolated components: optimization theory emphasizes training dynamics, generalization analysis concerns test performance, and strategies for initialization and deployment are often treated as independent stages. This fragmentation makes it impossible to characterize the coherent evolution of knowledge from prior injection, through data-driven updates, to task generalization. For example, the PAC-Bayes framework connects priors with generalization error but neglects the temporal causality of the training process; meta-learning involves multi-stage adaptation but fails to establish a symbolic model for state transitions. The absence of lifecycle modeling makes systematic analysis of the core question——“how parameters carry task knowledge from initialization through training”—difficult for model designers.
Second, there is a lack of explicit expression for the causal chain of information flow. Traditional approaches often treat the transformation from data to model as a black box, reducing the parameter update driven by loss function gradients to a purely mathematical optimization process and overlooking the causal dependencies among prior knowledge, data distribution, and model architecture. While the field of causal inference attempts to characterize causal relationships among variables, its application in machine learning is mostly limited to feature-label association analysis and fails to construct a complete causal chain from “prior constraints → initialization → gradient propagation → generalization capability.” For instance, adversarial training enhances robustness by maximizing loss, but theoretical analyses rarely explicitly link initial parameter constraints with the causal mechanisms underlying final adversarial robustness, making strategy design reliant on empirical trial-and-error.
Third, the mathematical foundations of interpretability tools are weak Ribeiro2016 ; Lundberg2017 ; Venkata2025 ; Lee2023 ; Alvarez2018 . Most current interpretability methods use local approximation or game-theoretic approaches to assign feature importance, which, while practically useful, lack rigorous mathematical axiomatization. For example, SHAP values compute marginal contributions in cooperative games but do not formalize their mapping to the information carrier states of models, potentially violating causality in interpretations. In addition, heuristic algorithms that infer data-feature dependencies from parameters lack mathematical guarantees for “information recoverability.” This theoretical weakness makes interpretability results susceptible to adversarial attacks and limits their applicability in high-stakes scenarios such as medicine or law.
Information science, as a core discipline connecting data, knowledge, and systems, offers a new perspective for the theoretical integration of machine learningShwartz2017 ; Xu2022 . Objective Information Theory (OIT) establishes a sextuple model of information through axiomatic methods, representing the intrinsic meaning of information via ontological states, reflecting its objective expression through carrier states, and explaining information realization through enabling mappings that possess both mathematical surjectivity and physical causality. This provides a universally applicable foundational model for understanding information in the objective world, especially within computer and information systemsXu2014 ; xu2015 ; Jianfeng2021 ; Jianfeng2023 ; Xu2024 ; Xu2025 .
On this basis, we rigorously define higher-order logic systems using recursion, and establish formal state expressions through interpretations of sets of well-formed formulas (WFFs), completing the last theoretical gap in OIT regarding the mathematical essence of state expressions. We then thoroughly apply the theory and methods of OIT to unify the core concepts throughout the machine learning lifecycle—prior knowledge encoding, data-driven training, and model generalization—as formal state expressions; we establish a mathematical logic system covering the entire ML process. Using this, we analyze and prove the causal-chain logic and constraint rules for typical stages in machine learning, including the formal modeling and realization of training information, the formal modeling and realization of model learning, and the model’s information processing and learning optimization, thus forming a Meta-Framework for Machine Learning Theory (MLT-MF). Finally, using MLT-MF, we propose universal definitions of model interpretability and ethical safety, prove and validate—via a simple feedforward neural network—three important theorems: model interpretability and information existence, ethical safety assurance, and an upper bound estimate on total variation distance (TVD). These results provide a powerful tool, based on a unified formal logic framework, for both theoretical analysis and engineering practice in machine learning.
2 Postulates and Formal Definition of Information
Undoubtedly, as a key component of artificial intelligence, modern machine learning is a product of the rapid development and high integration of information technologies such as computers, the Internet, big data, and cloud computing. Machine learning systems are, without question, an important branch of information systems. Starting from the basic concept of information can provide a unified theoretical foundation for research in machine learning and artificial intelligence.
2.1 Postulates of Information
Although humanity has entered the information age, the theoretical foundations of information science have long been dispersed across disciplines such as information theory, cybernetics, signal processing, statistical learning, and computational theory, each developing independently. There remains a lack of consensus on the concept of information, as well as a unified axiomatic basis and theoretical system. This is also a major reason why artificial intelligence often lacks interpretability and predictability.
To address this, we first propose the concept of an enabling mapping.
Definition 1 (Enabling Mapping) Let $X$ and $Y$ be two sets, and let $f$ be a mapping. If for any $x\in X$, there exists a $y\in Y$, and for any $y\in Y$, there exists at least one $x\in X$ such that $y=f(x)$, and furthermore, the existence of $y$ is physically realized only because of $x$, then $f$ is called an enabling mapping from $X$ to $Y$. This is denoted as:
$f:X\rightrightarrows Y$ (2.1)
or
$f(X)\underline{\triangleright}Y$ (2.2)
[…]
If $f(X)\underline{\triangleright}Y$, we also denote this as $f^{-1}(y)\underline{\triangleleft}x$ or $x\underline{\triangleright}f^{-1}(y)$, to maintain consistency in the directionality of enabling.
Based on the concept of enabling mapping, Xu2024 proposed four postulates of information:
Postulate 1 (Binary Attribute) Every piece of information comprises two components: the ontology, which represents the intrinsic essence of the information, and the carrier, which embodies the objective form of the information.
Postulate 2 (Existence Duration) The content and form of information have their own respective temporal existences.
Postulate 3 (State Representation) The states of the ontology and the carrier of information have their own respective representations.
Postulate 4 (Enabling Mapping) The state of the ontology can be mapped onto the state of the carrier via an enabling mapping.
These four postulates are concise, align with common conceptions of information, and are compatible with the basic concepts of major branches of information science. Their scientific significance is reflected in:
  • Unification: Providing a common foundation for long-dispersed theories in information science;
  • Axiomatization: Allowing the definition and reasoning of information to be axiomatized and formalized, akin to mathematics and physics;
  • Universality: Facilitating application to cutting-edge fields such as the Internet, big data, artificial intelligence, cognitive science, and cyber-physical systems.
The concept of "enabling mapping" is novel. It requires not only that a surjective mapping from the ontological state to the carrier state be established mathematically, but also that the existence of the ontological state renders the carrier state an objective reality, highlighting the dual character of information as both a mathematical mapping and a physical realizationShannon1948 ; Landauer1996 ; Wiener2019 .
2.2 Formal Representation of States
The postulates of information emphasize that the state of the ontology enables a mapping to the state of the carrier. Therefore, establishing a formal representation of both ontological and carrier states is particularly important.
Definition 2 (Higher-Order Formal Logical System): Suppose a formal system $\mathcal{L}$ is described using the following symbol set:
Variables $x_{1},x_{2},\ldots$;
Individual constants $a_{1},a_{2},\ldots$;
Functions $f_{1}^{1},f_{2}^{1},\ldots,f_{1}^{2},f_{2}^{2},\ldots,f_{1}^{3},f_{2}^{3},\ldots$;
Parentheses $(,)$;
Predicates $A_{1}^{1},A_{2}^{1},\ldots,A_{1}^{2},A_{2}^{2},\ldots,A_{1}^{3},A_{2}^{3},\ldots$;
Logical connectives "not" $\sim$, "implication" $\rightarrow$;
Quantifier "for all" $\forall$;
A term in $\mathcal{L}$ is defined as follows:
(1) Variables and individual constants are terms;
(2) If $f_{i}^{n}(n>0,i>0)$ is a function in $\mathcal{L}$, and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ that do not contain $f_{i}^{n}$ itself, then $f_{i}^{n}(t_{1},\ldots,t_{n})$ is also a term in $\mathcal{L}$;
(3) If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$, and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ that do not contain $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1},\ldots,t_{n})$ is a term in $\mathcal{L}$.
(4) The set of all terms in $\mathcal{L}$ is generated by (1), (2), and (3).
Based on this, we can define atomic formulas in $\mathcal{L}$:
If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$ and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ not containing $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1},\ldots,t_{n})$ is an atomic formula in $\mathcal{L}$.
We then define well-formed formulas (WFFs) in $\mathcal{L}$:
(5) Every atomic formula in $\mathcal{L}$ is a WFF;
(6) If $\mathcal{A}$ and $\mathcal{B}$ are WFFs in $\mathcal{L}$, then $\sim\mathcal{A}$, $\mathcal{A}\rightarrow\mathcal{B}$, $(\forall x_{i})\mathcal{A}$, $(\forall f_{i}^{n})\mathcal{A}$, and $(\forall A_{i}^{n})\mathcal{A}$ are all WFFs in $\mathcal{L}$, where $x_{i}$, $f_{i}^{n}$, $A_{i}^{n}$ are variables, functions, and predicates in $\mathcal{L}$, respectively, and no predicate or function symbol is applied, either directly or indirectly, to itself.
(7) The set of all WFFs in $\mathcal{L}$ is generated by (5) and (6).
The higher-order formal logical system $\mathcal{L}$ defined here, based on the flexibility of predicate logic, can describe various relationships and constraints among objects in a unified way. Through higher-order quantification support, the nested definition of complex data and behaviors, and it explicitly prohibits anypredicate or function symbol from applying to itself—directly or indirectly—to avoid self-referential paradoxes.
For simplicity, $\mathcal{L}$ does not define connectives such as "and" $\land$, "or" $\lor$, "if and only if" $\leftrightarrow$, or the quantifier "exists" $\exists$, and other commonly used logical symbols. However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of symbols and can make many formulas more conciseHamilton1988 . Therefore, they are introduced as defined symbols in subsequent discussions.
Definition 3 (Interpretation of a Formal System) An interpretation $E$ of a formal system $\mathcal{L}$ over a domain $D_{E}$ assigns specific values to constants, elements of the domain $D_{E}$ to variables, and turns function and predicate symbols into functions and predicates over that domain $D_{E}$, thereby yielding a set of WFFs each with a truth value.
Definition 3 assigns concrete values and meaning to a set of formulas in a formal system $\mathcal{L}$ through interpretation, thereby providing a fundamental approach for formally describing the definitional, attributive, structural, relational, and evolutionary state properties of objects as both the subjective and objective worlds.
Postulate 5 (Expression of Object State):The state set of an object collection over a specified time set is constituted by its attributes—such as properties, structures, values, and relations—as well as the evolution rules and logical combinations among these attributes, and satisfies the following conditions:
(1) Domain invariance: definitions of the object set and time set remain unchanged.
(2) Numerical determinacy: concrete numerical values determine values of certain properties and forms.
(3)Attribute expressibility: properties, forms, values, relations are expressible via functions and predicates in a formal system satisfying Definition 2.
(4) Rule expressibility: evolution rules among attributes are expressible in such a formal system using time variables and implication, and contain no implicit self-reference.
(5) Logical coherence: no logical contradictions exist among attributes and rules, nor among any logical conjunctions thereof.
Theorem 1 (Formal Expression of States) If the state set $S(X,T)$ of objects $X$ over times $T$ satisfies Postulate 5, then it can be represented as an interpretation of a formal system $\mathcal{L}$ over the domain $X\times T$, and it is logically coherent.
Proof: Since the state set $S(X,T)$ of the object set $X$ over the time set $T$ satisfies Postulate 5, we construct the corresponding formal system $\mathcal{L}$ as follows:
By (1), the definitions of the elements in $X$ and $T$ remain unchanged, ensuring determinacy. The variables in the formal system $\mathcal{L}$ are defined to correspond to each element of $X\times T$, thus making $X\times T$ the domain of discourse for $\mathcal{L}$. Consequently, $S(X,T)$ has a well-defined domain.
By (2), the constants in $\mathcal{L}$ are defined to represent all possible values in $S(X,T)$.
By (3), the functions and predicates in $\mathcal{L}$ represent the properties, forms, values, and relations of $X$ over $T$ as captured in $S(X,T)$. These properties can be expressed through functions and predicates in the formal system $\mathcal{L}$ that conform to Definition 1.
By (4), consider two ternary predicates $A_{1}^{3}$ and $A_{2}^{3}$.And then let $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))$ and $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2}))$ be two properties of $S(X,T)$ corresponding to objects $x_{1},x_{2}\in X$, times $t_{1},t_{2}\in T(t_{1}<t_{2})$, and binary functions $f_{1}(x_{1},t_{1}),f_{2}(x_{1},t_{1})$. Suppose all $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))$ satisfy the rule of evolving from $t_{1}$ to $t_{2}$ into some $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2}))$. The formula $(\forall A_{1}^{3}\forall f_{1}^{2}\forall x_{1}A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))\rightarrow\exists A_{2}^{3}\exists f_{2}^{2}\exists x_{2}A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2})))$ is also a well-formed formula in $\mathcal{L}$, with respect to the interpretation of $x_{1},x_{2},t_{1},t_{2}$, and is able to express the corresponding rules in $S(X,T)$. For other rules in $S(X,T)$, the related elements in the domain $X\times T$ can similarly be interpreted and expressed via corresponding WFFs in $\mathcal{L}$, with no implicit logical self-reference.
Therefore, $S(X,T)$ constitutes an interpretation $E$ of $\mathcal{L}$ over the domain $X\times T$. By (5), both itself and its logical closure are logically consistent. The theorem is thus proved.
Note that Theorem 1 stipulates that variables in $\mathcal{L}$ are elements of $X\times T$; thus, in formulas, object and time elements appear in pairs $(x,t)$. In practice, some attributes may be time-invariant; then simplified expressions may be used to describe such attributes without time.
Theorem 1 indicates that, by employing the higher-order formal system $\mathcal{L}$, it is possible to describe the states of virtually all things in the world. Both theoretical and empirical analyses qiu2025 further demonstrate that $\mathcal{L}$ is broadly applicable to the expression of state in disciplines such as mathematics, economics, ecology, computer and information systems, and natural language. This provides a unified state modeling theory for interdisciplinary research, which is of significant importance to fields such as artificial intelligence, knowledge engineering, and complex systems.
2.3 Finite Automaton States
Using the formal expression of states, we can study in depth the special and important object of finite automata.
Theorem 2 (State Representation of Finite Automata): For any finite automaton $M$, one can describe its input, output, and state-transition behavior over a related time set $T$ by the state set $S(M,T)$.
Proof: Let $M=(Q,R,U,\delta,\lambda)$ be a finite automaton, and $T=\{t_{i}|i=1,\ldots,n\}$ the (strictly increasing) time points at which $M$ performs input, output, or state-transition actions. By finite automaton theory mathsea2002 , $Q=\{q_{t_{i}}|t_{i}=1,\ldots,n-1\}$, $R=\{r_{t_{i}}|t_{i}=2,\ldots,n\}$, and $U=\{u_{t_{i}}|t_{i}=1,\ldots,n\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
$\delta:U\times Q\to U$ is the next-state function;
$\lambda:U\times Q\to R$ is the output function.
For each state $u_{t_{i}}\in U$, define the state predicate $State^{3}$ by
$\varphi_{U}(t_{i})=State^{3}(M,t_{i},u_{t_{i}})$ (2.3)
to assert that at time $t_{i}$, $M$ is in state $u_{t_{i}}$$i=1,\ldots,n$.
For each input $q_{t_{i}}\in Q$, define the input predicate $Input^{3}$ by
$\varphi_{I}(t_{i})=Input^{3}(M,t_{i},q_{t_{i}})$ (2.4)
to assert that at time $t_{i}$, $M$ receives input $q_{t_{i}}$$i=1,\ldots,n-1$.
For each output $r_{t_{i}}\in R$, define the output predicate $Output^{3}$ by
$\varphi_{O}(t_{i})=Output^{3}(M,t_{i},r_{t_{i}})$ (2.5)
to assert that at time $t_{i}$, $M$ produces output $r_{t_{i}}$$i=2,\ldots,n$.
The transition function $\delta$ is captured by the well-formed formula
$\varphi_{\delta}(t_{i})=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\delta(u_{t_{i}},q_{t_{i}})\land Eq^{2}(u_{t_{i+1}},\delta(u_{t_{i}},q_{t_{i}}))\land\varphi_{U}(t_{i+1})$ (2.6)
[…]
Similarly, the output function $\lambda$ is given by
$\varphi_{\lambda}(t_{i})=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\lambda(u_{t_{i}},q_{t_{i}})\land Eq^{2}(r_{t_{i+1}},\lambda(u_{t_{i}},q_{t_{i}}))\land$ (2.7)
$IsElementof^{2}(r_{t_{i+1}},R)\land\varphi_{O}(t_{i+1})$
[…]
Thus the state set
$S(M,T)=$ $\{\varphi_{ij}(t_{i}),\varphi_{I}(t_{j}),\varphi_{O}(t_{k}),\varphi_{S}(t_{j}),\varphi_{A}(t_{j}),$ (2.8)
$i=1,\ldots,n,j=1,\ldots,n-1,k=2,\ldots,n\}$
describes $M$’s inputs, outputs, and state-transitions over $T$. Here $Cn(X)$ denotes the logical closure of the set of well-formed formulas $X$. This completes the proof.
It follows that every finite automaton can be formalized by the state set given in Theorem 2, which moreover exhibits atuomatic transtion behavior. Hence we introduce :
Definition 4 (Finite Automaton State): If a state set $S(X,T)$ expresses a finite automaton, then $S(X,T)$ is called a finite automaton state.
All practical information systems, including computing systems, are finite automata in the mathematical sense. Therefore, the concept of finite automaton state is of great significance in the study of machine learning.
2.4 Definition of Information
As shown inXu2024 , the necessary and sufficient conditions for satisfying Postulates 1-4 are the existence of the following non-empty sets: an ontological set $o$, a set of times for the occurrence of the ontology $T_{h}$, a set of ontological states $S_{o}(o,T_{h})$, a set of objective carriers $c$, a set of times for the reflection of the carrier $T_{m}$, and a set of carrier reflection states $S_{c}(c,T_{m})$, such that $S_{o}(o,T_{h})$ enables a mapping to $S_{c}(c,T_{m})$. Denote this mapping as $I$, or more simply, $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, which is called the sextuple model of information.
This leads to the general definition of information:
Definition 5 (Mathematical Definition of Information): Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain. The ontology $o\in\mathbb{O}\cup\mathbb{S}$, time of occurrence $T_{h}\subseteq\mathbb{T}$, ontological state set $S_{o}(o,T_{h})$ satisfying Postulate 5, carrier $c\subseteq\mathbb{O}$ , reflection time $T_{m}\subseteq\mathbb{T}$, and carrier reflection set $S_{c}(c,T_{m})$ satisfying Postulate 5 are all non-empty sets. Then, information $I$ is the enabling mapping from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$.The set of all information $I$ is called the information space, denoted by $\mathfrak{I}$, and is one of the three major constituents of the objective world.
It is important to note that both $S_{o}(o,T_{h})$ and $S_{c}(c,T_{m})$ are state representations that satisfy Postulate 5. According to Theorem 1, each can be interpreted as a set of well-formed formulas within some formal system. However, the corresponding formal systems for the ontology and the carrier need not be identical. This distinction reflects one of the essential characteristics of information: its ability to establish connections between disparate entities in the world.
Definition 5 does not require that the ontological states of information be mapped one-to-one onto the carrier states. Of course, in practical scenarios, one-to-one mappings often do exist, which can simplify the analysis considerably.
Definition 6 (Recoverable Information): Let $I=(o,T_{h},S_{o},c,T_{m},S_{c})$ still denote a bijection from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$, i.e., $I$ is invertible. Thus, there exists an inverse mapping $I^{-1}$ such that $I^{-1}(S_{c}(c,T_{m}))\underline{\triangleleft}S_{o}(o,T_{h})$. In this case, we say that information $I$ is recoverable, and $S_{o}(o,T_{h})$ is the recoverable state set of information $I$.
In particular, owing to the importance of information in information systems, we propose:
Definition 7 (Information in an Information System): Let $I=(o,T_{h},S_{o},c,T_{m},S_{c})$ be a recoverable information. If $S_{o}(o,T_{h})$ is a finite automaton state, then $c$ is called an information system, and $I$ is called information in the information system $c$.
For recoverable information, we also use $I^{-1}(S_{c}(c,T_{m}))\underline{\triangleleft}S_{o}(o,T_{h})$ or $I^{-1}:S_{c}(c,T_{m})\leftleftarrows S_{o}(o,T_{h})$ to maintain consistency in the direction of enabling. Since, as recoverable information, the carrier losslessly carries the complete state of the ontology, this naturally represents an ideal scenario in information studies. If $I$ is not a bijection, we can always partition elements of $S_{o}(o,T_{h})$ according to the same image equivalence classes, and then establish an enabling mapping from these equivalence classes to $S_{c}(c,T_{m})$ by means of equivalence class reduction, thereby ensuring recoverability. Specifically:
Proposition 1 (Recoverable Reduction of Information): For $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, let $R$ be an equivalence relation on $S_{o}(o,T_{h})$ defined by $xRy\leftrightarrow I(x)=I(y)$. Then, there exists a unique enabling mapping $\hat{I}:S_{o}(o,T_{h})/R\rightrightarrows S_{c}(c,T_{m})$, and $\hat{I}$ is a bijection, i.e., $\hat{I}$ is recoverable. In this case, $\hat{I}$ is called the recoverable reduction of information $I$.
When $I$ is already recoverable, $S_{o}(o,T_{h})/R=S_{o}(o,T_{h})$, and $\hat{I}=I$. Recoverable information and the recoverable reduction of information can provide significant assistance in simplifying problems.
Reference Shannon1948 is generated as a foundational work in information theory. The paper proposed that $H=-\sum p_{i}\log p_{i}$ plays a central role as the measure of information in information theory." The discussion in this paper includes both the source and channel of information. The source is the origin of the information, and it generates events $e_{1},e_{2},\ldots,e_{n}$ with probabilities $p_{1},p_{2},\ldots,p_{n}$, respectively. The channel $I_{c}$ is the medium through which information is transmitted. After an event $e_{i}$ occurs at time $T_{h}$ in the source $I_{s}$, the channel $I_{c}$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$(where $1\leq i\leq n,\tau>0$).We define the ternary predicate $IsEvent^{3}(I_{s},T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $I_{s}$ at time $T_{h}$, and the ternary predicate $IsCode^{3}(I_{c},T_{m},c_{i})$ to represent that the channel $I_{c}$ received and transmitted code $c_{i}$ in time interval $T_{m}$. Then,
$S_{I_{s}}(I_{s},T_{h})=IsEvent^{3}(I_{s},T_{h},e_{i})$ (2.9)
and
$S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))=IsCode^{3}(I_{c},(T_{h},T_{h}+\tau),c_{i})$ (2.10)
are the states of the source $I_{s}$ at time $T_{b}$ and the state of the channel $I_{c}$ in the time interval $(T_{h},T_{h}+\tau)$, respectively (where $1\leq i\leq n)$. Furthermore,
$I:S_{I_{s}}(I_{s},T_{h})\rightrightarrows S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$ (2.11)
represents the functional enabling mapping from the source state $S_{I_{s}}(I_{s},T_{h})$ to the channel state $S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$.From equation (2.10), we can recover the event $e_{i}$ in equation (2.9) using the code $c_{i}$, thus $I$ is still recoverable information. Therefore, although Shannon’s information theory investigates the various probability distributions of $S_{I_{s}}(I_{s},T_{h})$ and the complex implementation mechanisms of $S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$, it forms a rich and extensive theoretical system. The OIT definition of information can encompass Shannon’s most original model of information, indicating that it has a broader and more universal application range.
2.5 Transmissibility of Information
Xu2014 ; xu2015 discuss transmissibility using composite mappings. Building on Definitions 1 and 5, we can characterize the fundamentals of information transmission more rigorously:
Theorem 3 (Transmissibility of Information) If $I_{1}=\langle o,T_{h},S_{o},c_{1},T_{m_{1}},S_{c_{1}}\rangle$ is information from an ontology $o$ to a carrier $c_{1}$, and $I_{2}=\langle c_{1},T_{m_{1}},S_{c_{1}},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is information from carrier $c_{1}$ to carrier $c_{2}$, then the composition $I_{3}=\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is also information, which we call transmission of information from $I_{1}$ via carrier $c_{1}$ to carrier $c_{2}$.
Proof: According to Definition 1, Definition 5, and the given conditions, for any $S_{o}(x,t)\in S_{o}(o,T_{h})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{1}(S_{o}(x,t))\underline{\triangleright}S_{c_{1}}(x_{1},t_{1})$. Furthermore, there exists $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$ such that $I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$. We define $I_{3}(S_{o}(x,t))=S_{c_{2}}(x_{2},t_{2})$. Since the physical realization of $S_{c_{2}}(x_{2},t_{2})$ depends on the existence of $S_{c_{1}}(x_{1},t_{1})$, and the existence of $S_{c_{1}}(x_{1},t_{1})$ in turn depends on $S_{o}(x,t)$, it follows by transitivity of causality that the realization of $S_{c_{2}}(x_{2},t_{2})$ ultimately depends on $S_{o}(x,t)$.
Conversely, for any $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$; and there also exists $S_{o}(x,t)\in S_{o}(o,T_{h})$ such that $I_{1}(S_{o}(x,t))\underline{\triangleright}S_{c_{1}}(x_{1},t_{1})$. Thus, $I_{3}(S_{o}(x,t))=I_{2}(I_{1}(S_{o}(x,t)))\underline{\triangleright}I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$.
Therefore, $I_{3}=\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is an enabling mapping from $S_{o}(o,T_{h})$ to $S_{c_{2}}(c_{2},T_{m_{2}})$, which satisfies the definition of information as given in Definition 5. In essence, this is the information $I_{1}$ transmitted from carrier $c_{1}$ to carrier $c_{2}$. The theorem is thus proved.
Transmission is an important way of information flow. Communication, processing, input, and output are all concrete manifestations of information transmission. In many cases, information undergoes more complex flows through chain-like transmission mechanisms. From Theorem 3, we can readily derive the following:
Corollary 1 (Information Chain) Let $\{I_{i}=\langle o,T_{h},S_{o},c_{i},T_{m_{i}},S_{c_{i}}\rangle|i=1,\ldots,n\}$ be a sequence of information mappings, where for any $i>1$, it holds that $c_{i-1}=o$, $T_{h}=T_{m_{i-1}}$, and $S_{o}(o,T_{h})=S_{c_{i-1}}(c_{i-1},T_{m_{i-1}})$. In this case, we say that these mappings constitute an information chain. Moreover, for any $1\leq i\leq j\leq n$, $I_{ij}=\langle o,T_{h},S_{o},c_{j},T_{m_{j}},S_{c_{j}}\rangle$ also qualifies as information according to Definition 5.
2.6 Noisy Information
Many researchers question the significance of recoverable information, as in the real world, information is often not perfectly recoverable. This is because the reflection of information is often influenced by noise or other external factors. Therefore, the following concept is introduced:
Definition 8 (Noisy Information): For information $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, let $\widetilde{S_{o}}(o,T_{h})$ be another state of the same carrier $c$ at time $T_{m}$ and $\widetilde{S_{c}}(c,T_{m})$ be another state of the same carrier $c$ at time $T_{m}$,
[…]
where the state $S^{-}_{o}(o,T_{h})$ is the loss state of $S_{o}(o,T_{h})$, satisfying $S^{-}_{o}(o,T_{h})\subseteq$ $S_{o}(o,T_{h})$ and not being enabled to map to $\widetilde{S_{c}}(c,T_{m})$, and the state $S^{+}_{o}(o,T_{h})$ is the superposed state of $S_{o}(o,T_{h})$, satisfying $S_{o}(o,T_{h})\cap S_{o}^{+}(o,T_{h})=\emptyset$ and being enabled to map to $\widetilde{S_{c}}(c,T_{m})$. The enabling mapping is then defined as:
$\tilde{I}:\widetilde{S_{o}}(o,T_{h})\rightrightarrows\widetilde{S_{c}}(c,T_{m})$ (2.13)
which is the noisy information of $I$, while $I$ is its own noiseless information. Thus, the noisy information $\tilde{I}$ is also information. It has the same ontology, time of occurrence, carrier, and reflection time as $I$, but the states of the entity and carrier may overlap but are not completely the same. Further analysis of equation (2.12) shows that information $I$ and $\tilde{I}$ are completely asymmetric, and hence:
Proposition 2 (Symmetry of Noisy Information): If $\tilde{I}$ is the noisy information of $I$, then $I$ is also the noisy information of $\tilde{I}$.
In fact, for both information $I$ and its noisy information $\tilde{I}$, it is easy to verify its recoverability or recoverable reducibility from the perspective of enabling mapping. However, in real scenarios, it is often difficult to obtain noise-free information $I$, and one can only study using a noisy version $\tilde{I}$. Moreover, the recoverable state of $\tilde{I}$ is not the same as that of $I$, which is the fundamental reason for the skepticism about the recoverability of information. But under the condition where the noiseless information IIcannot be directly obtained, we can approximate the ideal state of $I$ using $\tilde{I}$ along with the relationships between $S_{o}(o,T_{h}),S^{-}_{o}(o,T_{h})$, and $S_{o}^{+}(o,T_{h})$. This is the practical significance of introducing the concept of noisy information.
Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenarioShannon1948 . In this case, Shannonassumes that the signal generated by the source is $S$, the channel noise is $N$, and the channel output is:
$E=f(S,N)$ (2.14)
where $f$ is the distortion function of the input for the transmitted signal. According to OIT’s view, the signal $S$ here is the ontological state of information $I$, $S_{I_{s}}(I_{s},T_{h})$. The distortion function $f(S,N)$ for signal $S$ and noise $N$ is equivalent to removing a state $S_{I_{s}}^{-}(I_{s},T_{h})$ from the source state $S_{I_{s}}(I_{s},T_{h})$ and adding another state $S_{I_{s}}^{+}(I_{s},T_{h})$. The output $E$ is then the channel’s output state $\widetilde{S_{I_{c}}}(I_{c},(T_{h},T_{h}+\tau))$, where $(T_{h},T_{h}+\tau)$ is the time interval of the channel output. Thus,
$\tilde{I}:\widetilde{S_{I_{s}}}(I_{s},T_{h})\rightrightarrows\widetilde{S_{I_{c}}}(I_{c},(T_{h},T_{h}+\tau))$ (2.15)
is the noisy information of
$I:S_{I_{s}}(I_{s},T_{h})\rightrightarrows S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$ (2.16)
[…]
In this way, we use the definition of noisy information to explain Shannon’s model of noisy channels inShannon1948 , demonstrating that OIT is entirely applicable to the complex scenarios of information research under noisy conditions.
3 Meta-Framework for Machine Learning Theory (MLT-MF)
The core of machine learning is the artificial intelligence model, which can be regarded as a finite automaton as well as an information system. The processes of operation, learning, testing, and application for such a model all fundamentally depend on information within the system. Therefore, by employing the concepts of finite automaton state and information models established in Section 2, the process of machine learning and its applications can be unified into five primary steps of information realization and functioning:
3.1 Construction of Initial Information
Suppose the model under study is denoted as M, and let $T=\{t_{i}|i=1,\ldots,n\}$ be the set of time points at which M, as a finite automaton, undergoes assignment, input, output, and state transitions in increasing order. According to the perspective of OIT, the creation of M is essentially the realization of information
$I_{o}=\langle M_{D},T_{o_{h}},S_{M_{D}},M,t_{1},S_{M}\rangle$ (3.1)
where the ontology $M_{D}$ of $I_{o}$ is the designer’s conceptualization of M; $T_{o_{h}}$ is the time at which the designer conceives M, and $S_{M_{D}}(M_{D},T_{o_{h}})$ is the initial finite automaton state that must be realized at this conceptual stage. The carrier M of $I_{o}$ is the model itself; the reflection time $t_{1}$ is when M is constructed according to $S_{M_{D}}(M_{D},T_{o_{h}})$; and $S_{M}(M,t_{1})$ is the formalization of the actual physical state of M upon its completion. Thus, $\sup(T_{o_{h}})<\sup(t_{1})$.
According to Theorem 2, let the binary function $\Phi^{2}(X,t)$ represent the set of structures, parameters, and algorithms of model $X$ at time $t$. The formula
$\varphi_{o_{1}}=State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))$ (3.2)
denotes that, at time $t_{1}$, the initial structure, parameters, and algorithm set $\Phi^{2}(M_{D},T_{o_{h}})$ as envisioned by the designer are assigned to $M_{D}$.
$\varphi_{o_{2}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))\land Input^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
$Learnable^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$
$(Eq^{3}(\Phi^{2}(M_{D},t_{2}),f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))))\land State^{3}(M_{D},t_{2},\Phi^{2}(M_{D},t_{2}))$ (3.3)
expresses that, at time $t_{1}$, when the structure, parameters, and algorithm set of $M_{D}$ is $\Phi^{2}(M_{D},T_{o_{h}})$, and the input carrier state $S_{M_{I}}(M_{I},t_{1})$ is provided and satisfies the learnability condition determined by the binary predicate $Learnable^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$, $M_{D}$ can optimize its internal structures, parameters, and algorithmsto $\Phi^{2}(M_{D},t_{2})$ through the learning binary function $f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$, assigning this to $M_{D}$. This demonstrates the model’s capacity for learning and optimization, where $M_{I}$ denotes the input device of M.
$\varphi_{o_{3}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))\land Input^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
$Processable^{2}(M,S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1}))$
$(Eq^{3}(S_{M_{O}}(M_{O},t_{2}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1})))\land Output^{3}(M,t_{2},S_{M_{O}}(M_{O},t_{2})))$ (3.4)
indicates that, at time $t_{1}$, if $S_{M_{I}}(M_{I},t_{1})$ is input into $M_{D}$ and meets the processability condition specified by the binary predicate $Processable^{2}(M,S_{M_{I}}(M_{I},t_{1}))$, then $M_{D}$ can process $S_{M_{I}}(M_{I},t_{1})$ via the binary processing function $f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1}))$ to yield the output carrier state $S_{M_{O}}(M_{O},t_{2})$, which is then output at time $t_{2}$. This demonstrates the model’s ability to process and output information, where $M_{O}$ denotes the output device of M.
Thus,
$S_{M_{D}}(M_{D},T_{o_{h}})=\{\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}\}$ (3.5)
constitutes the initial state of M as conceived by the designer.
It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},T_{o_{h}})$, and $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))$ all represent sets, $\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}$ may be considered as three well-formed formulas, or—when decomposed with respect to the elements of each set—as three sets of more granular well-formed formulas. Regardless of the situation, both cases satisfy Definition 4 concerning finite automaton states. Therefore, subsequent discussions will encompass both cases without further elaboration.
$S_{M_{D}}(M_{D},T_{o_{h}})$ may represent the designer’s idea, a design draft on paper, or an electronic document. Through the information $I_{O}$, these are enabled and mapped into the physically operational state of the model M:
$S_{M}(M,t_{1})=\{\psi_{o_{1}},\psi_{o_{2}},\psi_{o_{3}}\}$ (3.6)
where $\psi_{o_{1}}\underline{\triangleleft}I_{o}(\varphi_{o_{1}})$, $\psi_{o_{2}}\underline{\triangleleft}I_{o}(\varphi_{o_{2}})$, and $\psi_{o_{3}}\underline{\triangleleft}I_{o}(\varphi_{o_{3}})$ are the enabling mappings of $\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}$, respectively. This process establishes M as a typical finite automaton and an operational information system.
3.2 Formal Modeling and Realization of Training Information
According to the general formal definition of information, the formal expression for the realization of training information is straightforward:
$I_{t}=\langle ot,T_{th},S_{ot},M_{I},t_{1},S_{M_{I}}\rangle$ (3.7)
$S_{ot}(ot,T_{th})=\{\varphi_{t}|\varphi_{t}\text{ is a WFF about }ot\text{ and }T_{th}\}$ (3.8)
$S_{M_{I}}(M_{I},t_{1})=\{\psi_{t}|\psi_{t}\text{ is a WFF about }M_{I}\text{ and }t_{1}\}$ (3.9)
Here, the ontology $ot$ refers to the object containing knowledge or samples; $T_{th}$ is the time at which its state is formed; and $S_{ot}(ot,T_{th})$ is a set of business-oriented semantic assertions or facts about the ontology $M_{I}$ functions similarly to the hardware interface and API software of M, carrying the training information at time $t_{1}$. The carrier state $S_{M_{I}}(M_{I},t_{1})$ is the physical realization of $S_{ot}(ot,T_{th})$ via an enabling mapping, ensuring that the information can be machine-read and entered into M.
3.3 Formal Modeling and Implementation of Model Learning
Model learning for M involves fusing its original information $I_{o}$ with the training information $I_{t}$, resulting in updated information:
$I_{u}=\langle M\cup M_{I},t_{1}\cup t_{2},S_{M\cup M_{I}},M,t_{2},S_{M}\rangle$ (3.10)
Here, the ontology is the hybrid $M\cup M_{I}$, encompassing both the model and its interface; $t_{1}\cup t_{2}$ marks the time of the first learning event; and $S_{M\cup M_{I}}(M\cup M_{I},t_{1}\cup t_{2})$ formally represents the hybrid state during learning:
$S_{M\cup M_{I}}(M\cup M_{I},t_{1}\cup t_{2})=\{\varphi_{u_{1}},\varphi_{u_{2}},\varphi_{u_{3}}\}$ (3.11)
$\varphi_{u_{1}}=Eq^{3}(\Phi^{2}(M,t_{2}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{1})))\land State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))$ (3.12)
indicates that the input state $S_{M_{I}}(M_{I},t_{1})$, which was learned by M at time $t_{1}$, has been transformed into a new set of structures, parameters, and algorithms $\Phi^{2}(M,t_{2})$ at time $t_{2}$, thus becoming the state of M at $t_{2}$.
$\varphi_{u_{2}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
$Learnable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
$(Eq^{3}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land State^{3}(M,t_{3},\Phi^{2}(M,t_{3})))$ (3.13)
shows that after the first learning step, M remains capable of further learning and optimization.
$\varphi_{u_{3}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
$Processable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
$(Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land Output^{3}(M,t_{3},S_{M_{O}}(M_{O},t_{3})))$ (3.14)
indicates that when the training information $I_{t}$ is processable by M, it can also produce output information.
This state is then enabled and mapped to the physical state of the carrier M at time $t_{2}$:
$S_{M}(M,t_{2})=\{\psi_{u_{1}},\psi_{u_{2}},\psi_{u_{3}}\}$ (3.15)
3.4 Information Processing and Learning Optimization in Models
User input information is formally given by:
$I_{q}=\langle oq,T_{qh},S_{oq},M_{I},t_{2},S_{M_{I}}\rangle$ (3.16)
$S_{oq}(oq,T_{qh})=\{\varphi_{q}|\varphi_{q}\text{ is a WFF about }oq\text{ and }T_{qh}\}$ (3.17)
$S_{M_{I}}(M_{I},t_{2})=\{\psi_{q}|\psi_{q}\text{ is a WFF about }M_{I}\text{ and }t_{2}\}$ (3.18)
Here, the ontology $oq$ represents the user’s intention or device, with $T_{qh}$ denoting the time at which the ontological state is formed. The state $S_{oq}(oq,T_{qh})$ is the formal representation of the user’s needs. The carrier state $S_{M_{I}}(M_{I},t_{2})$ is the physical realization, via enabling mapping, of $S_{oq}(oq,T_{qh})$ onto the model’s interface device $M_{I}$.
3.5 Information Processing, Output, and Online Learning of the Model
When the model M processes input and produces output, the information it carries can be represented as:
$I_{v}=\langle M\cup M_{I}\cup M_{O},t_{2}\cup t_{3},S_{M\cup M_{I}\cup M_{O}},M,t_{3},S_{M}\rangle$ (3.19)
Where:
$S_{M\cup M_{I}\cup M_{O}}(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3})=\{\varphi_{v_{1}},\varphi_{v_{2}},\varphi_{v_{3}}\}$ (3.20)
$\varphi_{v_{1}}=State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))$ (3.21)
indicates that $\Phi^{2}(M,t_{2})$ has been assigned to M at time $t_{2}$.
$\varphi_{v_{2}}=$ $Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land$
$Output^{3}(M,t_{3},S_{M_{O}}(M_{O},t_{3})))$ (3.22)
indicates that when the input state $S_{M_{I}}(M_{I},t_{2})$ is received at $t_{2}$, M produces the output state $S_{M_{O}}(M_{O},t_{3})$ at $t_{3}$.
$\varphi_{v_{3}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
$Learnable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
$(Eq^{3}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land State^{3}(M,t_{3},\Phi^{2}(M,t_{3})))$ (3.23)
expresses that if the input information can be learned by M, then M can achieve online learning while processing and outputting information.
The physical realization of this ontological state through an enabling mapping is:
$S_{M}(M,t_{3})=\{\psi_{v_{1}},\psi_{v_{2}},\psi_{v_{3}}\}$ (3.24)
The output information is given by:
$I_{r}=\langle M\cup M_{I},t_{2},S_{M\cup M_{I}},M_{O},t_{3},S_{M_{O}}\rangle$ (3.25)
where the ontological state is:
$S_{M\cup M_{I}}(M\cup M_{I},t_{2})=\{\varphi_{r_{1}}\}$ (3.26)
with
$\varphi_{r_{1}}=Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))$ (3.27)
The formal expression of the physical state of the model’s output device $M_{O}$ at time $t_{3}$ is:
$S_{M_{O}}(M_{O},t_{3})=\{\psi_{r_{1}}\}$ (3.28)
The model can repeat the processes described in Sections 3.2 to 3.5 , iterating through cycles of training and information processing to meet user demands and continuously improve its performance.
3.6 Multi-Stage Information Causal Chains
From a systems engineering perspective, the realization of information in machine learning involves more than just five information mappings. In practice, stages such as requirements analysis, design and development, and system integration are also necessary. Specifically, in the four stages of model construction, learning and training, problem input, and answer output, and according to the corollary on information chains in Section 2.5, there is a causal and progressive process that moves from conceptual ideas to mind maps, program code, and finally to physical implementation. In this process, the earlier content serves as the ontological state, while the subsequent content becomes the carrier state, sequentially forming conceptual information, design information, and implementation information (see Figure.1).
Figure 1: Machine Learning Causality Chains with Multiple Information Mappings
As shown in Figure 1, each of these four stages encompasses four layers: conceptual ideas, mind maps, program code, and physical models. Typically, content within the same layer shares the same semantic logic and can thus be uniformly represented by a formal system. Furthermore, due to the transmissive nature of information, the innermost physical model can also be regarded as the result of enabling mappings from outer conceptual ideas or mind maps. This illustrates that the entire system comprises multiple stages, with each stage involving various combinations of information causal chains.
This framework, based on the realization of information and finite automaton states, provides a formalized representation and a comprehensive, causally progressive structure for each phase in the machine learning process. It transcends specific algorithms or models, offering a unified and universal framework, MLT-MF. In the appendix, we also present an example of its application to feedforward neural networks, which validates the correctness and applicability of the framework and supports the verification of the related theorems.
4 Several Important Theorems and Proofs
The MLT-MF framework is applicable to a range of key issues in the field of artificial intelligence, offering many unique and practical theoretical principles and methodological guidelines.
4.1 Theorem of Model Interpretability
Interpretability of complex models, such as those in deep learning, has long been one of the key challenges in artificial intelligence. However, the academic community has yet to establish a rigorous mathematical formulation of model interpretability Lipton2018 ; Doshi2017 ; Rubin2019 ; Miller2019 . In fact, interpretability should be understood as a concept that depends on the domain of discourse and the corresponding logical system. Furthermore, it should also be considered in relation to the specific inputs and outputs of the model. Based on MLT-MF, we define the following concepts:
Definition 9 (Model Interpretability) Let the model M, as a finite automaton, operate over an ordered set of times $T=\{t_{i}|i=1,\ldots,n\}$, at which it undergoes assignment, input, output, and state transitions. The state of M at time $t_{i}$ is denoted as $S_{M}(M,t_{i})$, and the state of its input device $M_{I}$ at time $t_{i}$ is $S_{M_{I}}(M_{I},t_{i})$, while the state of the output device $M_{O}$ at time $t_{i+1}$ is $S_{M_{O}}(M_{O},t_{i+1})$. If there exists a formal system $\mathcal{L}$ along with three interpretation relations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$ respectively, such that
$E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$ (4.1)
making $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ the respective enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, then M is said to be interpretable in $\mathcal{L}$ with respect to output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$. If for all $i=1,\cdots,n-1$, the output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in $\mathcal{L}$, then M is called interpretable in $\mathcal{L}$.
Here, $\text{Cn}(E_{M}\cup E_{M_{I}})$ denotes the logical closure of $E_{M}\cup E_{M_{I}}$. Equation 4.1 states that in the formal system $\mathcal{L}$, all formulas in $E_{M_{O}}$ are logical consequences of $E_{M}$ and $E_{M_{I}}$. Moreover, since $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are respectively enabled by $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, the formal system $\mathcal{L}$ and its three interpretations fully capture the logical and causal relationship whereby M produces output $S_{M_{O}}(M_{O},t_{i+1})$ under input $S_{M_{I}}(M_{I},t_{i})$. This, in essence, offers a concise understanding of model interpretability.
Theorem 4 (Model Interpretability and Information Existence) Let the model M be viewed as a finite automaton operating over an ordered set of times $T=\{t_{i}|i=1,\ldots,n\}$, during which it undergoes assignment, input, output, and state transitions.
(1) For any $1\leq i\leq n-1$, a necessary and sufficient condition for the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M to be interpretable in the formal system $\mathcal{L}$ under the input $S_{M_{I}}(M_{I},t_{i})$ is that there exist $Iu(t_{i})=\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$ and $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$ representing the information carried by M and its input device $M_{I}$ at time $t_{i}$ respectively, and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$ representing the information carried by the output device $M_{O}$ at time $t_{i+1}$. The ontological states of these three pieces of information must each correspond to an interpretation in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$ must hold.
(2) A necessary and sufficient condition for M to be interpretable in $\mathcal{L}$ is that, for all $i=1,\cdots,n-1$, there exist information objects $Iu(t_{i})$, $Iq(t_{i})$, and $Ir(t_{i+1})$ whose ontological states are all interpretations in $\mathcal{L}$, and such that the ontological state of $Ir(t_{i+1})$ is contained in the logical closure of the union of the ontological states of $Iu(t_{i})$ and $Iq(t_{i})$.
Proof: For any $1\leq i\leq n-1$, suppose that the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M under the input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in the formal system $\mathcal{L}$. According to Definition 9, there exist three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ in the formal system $\mathcal{L}$, respectively over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$, such that $E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$, and $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are each the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
Let $S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))=E_{M}$, $S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))=E_{M_{I}}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))=E_{M_{O}}$. Then, according to Definition 6,we find there exist information objects $Iu(t_{i})=$ $\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$, similarly , then $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$, and $Ir(t_{i+1})=\langle or(t_{i+1}),$ $Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$, such that $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$. Thus, the necessity of the first part of the theorem is established.
On the other hand, for any $1\leq i\leq n-1$, suppose there exist some information objects $Iu(t_{i})=\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$, $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$, and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$, such that the ontological states $S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))$, $S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))$ are interpretations in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$. Then, according to Theorem 1, $E_{M}=S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times Tu_{h}(t_{i})$, $oq(t_{i})\times Tq_{h}(t_{i})$, and $or(t_{i+1})\times Tr_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$. At the same time, $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively. By Definition 9, the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M under the input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in the formal system $\mathcal{L}$. Therefore, the sufficiency of the first part of the theorem is established.
For all $i=1,\ldots,n-1$ , by combining Definition 9 with the first part of the theorem, the second part of the theorem follows directly. This completes the proof.
4.2 Theorem of Ethical Safety Assurance
Ethical safety has long been one of the most widely discussed and scrutinized issues in the field of artificial intelligence mathsea2002 ; Doshi2017 ; Rubin2019 ; Miller2019 , and to date, there is still no universally accepted and practically actionable rigorous definition Russell2015 ; Jobin2019 ; Corrêa2023 . Based on the MLT-MF framework, we are able to theoretically ensure the ethical safety of artificial intelligence through the definition of ethical safety and the corresponding formal expressions. To achieve this, it is first necessary to specify appropriate well-formed formulas $Ec$ as ethical constraints—such as those governing fairness and privacy protection—thus establishing a formalized definition of ethical safety.
Definition 10 (Ethical Safety): Let $S(X,T)$ be the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$. Let $Ec(\vec{v})$ denote an ethical constraint expressed as a well-formed formula in $\mathcal{L}$, where $\vec{v}=(v_{1},\ldots,v_{n})$ are its free variables. For each substitutable $(x,t)\in X\times T$, define
$Ec|_{(x,t)}=Ec(\vec{v}\mapsto\vec{\sigma}(x,t))$ (4.2)
where $\vec{\sigma}:X\times T\to D^{n}$ is a value vector function such that $\sigma_{i}(x,t)\in D$ is the concrete value assigned to $v_{i}$ for $1\leq i\leq n$.
If for all $\varphi\in\text{Cn}(S(X,T))$, and all $(x,t)\in X\times T$,
$\varphi\rightarrow\sim Ec|_{(x,t)}$ (4.3)
does not hold, then $S(X,T)$ is said to be ethically safe with respect to $Ec$. Furthermore, if a model M is capable of producing output $I$, and both the ontological and carrier states of $I$ are ethically safe with respect to $Ec$, then M is said to be ethically safe with respect to $Ec$.
In the definition of ethical safety, we require not only that $S(X,T)$ contains no formulas that entail $\sim Ec|_{(x,t)}$, but also that its logical closure $\text{Cn}(S(X,T))$ contains no such formulas. This is because if $\text{Cn}(S(X,T))$ were to include formulas entailing $\sim Ec|_{(x,t)}$, then the conjunction of multiple formulas in $S(X,T)$ could still violate the ethical constraint $Ec$. This would indicate that $S(X,T)$ implicitly allows situations that contravene $Ec$, which must be strictly avoided in a rigorous sense. Moreover, it is generally not difficult for a model M to achieve ethical safety with respect to $Ec$ in a simple manner; the key challenge lies in ensuring the largest possible set of valid outputs while maintaining ethical safety. To this end, we propose:
Theorem 5 (Ethical Safety Assurance): Suppose model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$. Let $S_{or}(or,Tr_{h})$ denote the ontological state of $Ir$. For any ethical constraint $Ec$ represented by a well-formed formula in $\mathcal{L}$, there exists a constructive algorithm for obtaining the largest output state subset $S_{or}^{EC}(or,Tr_{h})\subseteq S_{or}(or,Tr_{h})$ that is ethically safe with respect to $Ec$. By introducing an ethical safeguard formula for $Ec$ into $Iu$, it is thereby possible to ensure that model M attains the maximal output state that is ethically safe with respect to $Ec$.
Proof: Since model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$, for the ontological states $S_{ou}(ou,Tu_{h})$, $S_{oq}(oq,Tq_{h})$ and $S_{or}(or,Tr_{h})$ of $Iu$, IqIq and $Ir$ respectively, the MLT-MF framework provides:
$\varphi u=($ $\forall\Phi u)(\forall\Phi q)(\text{IsSubsetOf}^{2}(\Phi u,Sou(ou,Tu_{h}))\land\text{IsSubsetOf}^{2}(\Phi q,S_{oq}(oq,Tq_{h}))\land$
$\text{Processable}^{2}(\Phi u,\Phi q))\to\exists\varphi r(\text{IsFormulaOf}^{2}(\varphi r,S_{or}(or,Tr_{h}))\land$
$\text{Eq}^{2}(\varphi r,f_{process}^{2}(\Phi u,\Phi q)))$ (4.4)
Thus, the set
$S_{or}(or,Tr_{h})=\{\varphi r|\varphi r=\text{Processable}^{2}(\Phi u,\Phi q),\Phi u\in Sou(ou,Tu_{h}),\Phi q\in S_{oq}(oq,Tq_{h})\}$ (4.5)
represents the theoretical output states of model M.
Since $S_{or}(or,Tr_{h})$ is a finite set, we construct a finite hypergraph
$H_{or}=(S_{or}(or,Tr_{h}),\mathcal{E})$ (4.6)
where the vertex set is precisely $S_{or}(or,Tr_{h})$, and the hyperedge set $\mathcal{E}$ covers all well-formed formulas that directly or implicitly violate the ethical constraint $Ec$:
(1) For any $\varphi r\in S_{or}(or,Tr_{h})$ and $(x,t)\in or\times Tr_{h}$, if $\varphi r\rightarrow\sim Ec|_{(x,t)}$, then add the hyperedge $\{\varphi r\}$ to $\mathcal{E}$;
(2) If there exists a minimal subset $W\subseteq S_{or}(or,Tr_{h})$ such that the conjunction of all formulas in $W$ entails $\sim Ec|_{(x,t)}$, then add the hyperedge $W$ to $\mathcal{E}$.
Let $G$ be an independent set in $H_{or}$, i.e., a subset of vertices containing no hyperedge $e\in\mathcal{E}$. If there exists $\varphi\in\text{Cn}(G)$ such that
$\varphi\to\sim Ec|_{(x,t)}$ (4.7)
then there must exist a finite subset $\{\varphi r_{1},\cdots,\varphi r_{k}\}\subseteq G$ such that $\varphi r_{1}\land\cdots\land\varphi r_{k}=\varphi$, and hence $\varphi r_{1}\land\cdots\land\varphi r_{k}\to\sim Ec|_{(x,t)}$. Let $W=\{\varphi r_{1},\cdots,\varphi r_{k}\}$, so the conjunction of $W$ entails $\sim Ec|_{(x,t)}$.
According to the hyperedge construction rules, if $|W|=1$, then $W$ itself directly violates $Ec$, and there must be a hyperedge $\{\varphi r_{1}\}$. But since $\{\varphi r_{1}\}\in G$, this contradicts the assumption that $G$ is an independent set. If $|W|\geq 2$, then by the compactness theorem, the logical entailment can be captured by a finite minimal subset, so there exists a minimal subset $W_{min}\subseteq W$ whose conjunction entails $\sim Ec|_{(x,t)}$. Therefore, there exists a hyperedge $W_{min}\subseteq W\subseteq G$ in $H_{or}$, again contradicting the independence of $G$. Thus, (4.7) cannot hold, and the independent set $G$ is ethically safe with respect to $Ec$.
By the maximum independent set theorem for finite hypergraphs, $H_{or}$ necessarily contains an independent set $G_{max}$ of maximal cardinality. Suppose there exists another ethically safe subset $K$ with
$|K|>|G_{max}|$ (4.8)
then $K$ cannot contain any hyperedge $e\in{E}$. Otherwise, if $|e|=1$, then $K$ directly violates $Ec$; if $|e|\geq 2$, then the formulas in $K$ violate $Ec$. Both cases contradict the ethical safety assumption for $K$. Therefore, $K$ must alsobe an independent set in $H_{or}$, which contradicts the maximality of $G_{max}$. Hence, $G_{max}$ is the largest ethically safe subset of $S_{or}(or,Tr_{h})$ with respect to $Ec$.
[…]
In this way, add to $Sou(ou,Tu_{h})$ an ethical safeguard formula:
$\varphi u=($ $\forall\varphi r)(IsFormulaOf^{2}(\varphi r,S_{or}(or,Tr_{h})))\to\exists\varphi s((IsFormulaOf^{2}(\varphi s,S_{os}(os,Ts_{h}))\land$
$(IsFormulaOf^{2}(\varphi r,S_{or}^{Ec}(or,Tr_{h}))\to Eq^{2}(\varphi s,\varphi r))\land$
$(IsFormulaOf^{2}(\varphi r,S_{or}^{\sim Ec}(or,Tr_{h}))\to Eq^{2}(\varphi s,IsSafeback^{1}(NEc))))$ (4.9)
Since the newly added $\varphi u$ pertains to $S_{os}(os,Ts_{h})$ and is entirely independent of the existing formulas in $Sou(ou,Tu_{h})$, it does not affect the satisfaction of Postulate 5 by $Sou(ou,Tu_{h})$. Moreover,
$S_{os}(os,Ts_{h})=S_{or}^{Ec}(or,Tr_{h})\cup\{IsSafeback^{1}(NEc)\}$ (4.10)
Because NEc and IsSafeback1(NEc) were not previously present in $\mathcal{L}$, and $S_{or}^{Ec}(or,Tr_{h})$ and $\{IsSafeback^{1}(NEc)\}$ are logically independent and both logically consistent, $S_{os}(os,Ts_{h})$ is also logically consistent and satisfies all other conditions of Postulate 5. Therefore, it is the maximal output state that is ethically safe with respect to $Ec$.
By taking $S_{os}(os,Ts_{h})$ as the ontological state and mapping it via recoverable information $Is$ to the carrier state $S_{M}(M,Ts_{m})$, $S_{M}(M,Ts_{m})$ is necessarily also ethically safe with respect to $Ec$. This satisfies the condition in Definition 10 for model M to be ethically safe with respect to $Ec$. The theorem is thus proved.
Theorem 5 not only establishes that it is possible to construct an ethical safety assurance mechanism for models, but also provides a constructive algorithm—based on hypergraphs-for obtaining the maximum ethically safe output. However, computing the maximum independent set in a hypergraph is generally an NP-hard problem. In practical applications, the algorithm can be simplified by leveraging two functions: one for ranking the conflict intensity among formulas in $S_{or}(or,Tr_{h})$, and another for ethical safety checking, thus yielding an approximately maximal safe output. This is a practical issue that warrants further investigation and will not be elaborated upon here. Meanwhile, the inheritance of ethical safety in the learning process can be achieved by incorporating dedicated ethical constraint regularization terms into the model algorithm. This involves more specific model structures and algorithmic design, which are beyond the scope of this paper.
In summary, based on the MLT-MF framework, Definition 10 and Theorem 5, for the first time, construct a formal theoretical framework for ethical safety. This approach, unlike those based on heuristic rules or statistical indicators, fills the gap in mathematical foundations and verifiability in traditional research, and provides concrete methodological guidance for the automated embedding and verification of ethical constraints.
4.3 Theorem on Model Generalization Error Estimation
Improving model generalization ability is among the most critical objectives in machine learning. However, in practice, generalization error is usually difficult to analyze and derive theoretically in advance Vapnik1999 ; Bartlett2002 ; Zhang2016 ; Advani2020 ; Perlaza2024 . It can typically only be estimated through testing after training, and it is challenging to obtain accurate error metrics from limited test samples.Applying the MLT-MF framework allows us to estimate upper bounds on the generalization error measured by Total Variation Distance (TVD) between model distributions, facilitating the analysis and prediction of model generalization performance across a variety of tasks.
Theorem 6 (Upper Bound on TVD) Let the information describing the model M after learning and optimization be $I_{u}=\langle ou,Tu_{h},S_{ou},M,Tu_{m},S_{M}\rangle$ and the input information be $I_{q}=\langle oq,Tq_{h},S_{oq},eq,Tq_{m},S_{cq}\rangle$. Assume both $S_{ou}$ and $S_{oq}$ are distributed according to probability distributions over the same measurable space. Let $|S_{oq}|$ and $|S_{ou}\cap S_{oq}|$ denote the cardinalities of the finite sets $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively. The probability distribution of the model is $p_{M}(x)$, where $p_{M}(x)>0$ for $x\in S_{ou}$, the distribution over $S_{oq}$ is $p_{q}(x)$,with $p_{q}(x)>0$ for $x\in S_{oq}$ . The total probability mass of $p_{q}(x)$ over $S_{ou}\cap S_{oq}$ is denoted $P_{Mq}$. When the model M is sufficiently trained on $S_{ou}$ and adopts a conservative uniform prediction for unknown states, the total variation distance between $p_{M}(x)$ and $p_{q}(x)$ satisfies:
$TVD(p_{q},p_{M})\begin{cases}\approx 0,&S_{oq}\subseteq S_{ou}\\ \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log(\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}})-H(p_{q}))},&S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (4.11)
For the second case, it is assumed that $S_{oq}\setminus S_{ou}\neq\emptyset$, $1-P_{Mq}>0$, and $|S_{oq}|-|S_{ou}\cap S_{oq}|>0$. Here, $H(p_{q})=-\sum_{x\in S_{oq}\setminus S_{ou}}p_{q}(x)\log p_{q}(x)$ is the entropy of the data in the non-overlapping region.
[…]
$p_{M}(x)\begin{cases}\approx p_{q}(x),&x\in S_{ou}\cap S_{oq}\\ =(1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|),&x\in S_{oq}\backslash S_{ou}\end{cases}$ (4.12)
Using KL divergence to measure the discrepancy between the true distribution and the hypothesized model distribution, we have Goodfellow2016 :
$D_{KL}(p_{q}(x)\|p_{M}(x))=$ $\sum_{x\in S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ (4.13)
$=$ $\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$
According to (4.12), in the overlapping region:
$\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}\to\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{q}(x)}=0$ (4.14)
This establishes the first case in (4.11).
In the non-overlapping region, the true distribution $p_{q}(x)$ can be arbitrary. Similarly, by (4.12), we have
$\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ $=\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{q}(x)-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log p_{M}(x)$
$=-H(p_{q})-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))$
$=-H(p_{q})-\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)$ (1)
$=(1-P_{Mq})\log((|S_{oq}|-|S_{ou}\cap S_{oq}|)/(1-P_{Mq}))-H(p_{q})$ (4.15)
Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their total variation distance. By Pinsker’s inequality Pinsker1964 , we have:
$\text{TVD}(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\|p_{M}(x))}$
$=\sqrt{\frac{1}{2}(\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)})}$
$\approx\sqrt{\frac{1}{2}((1-P_{Mq})\log((|S_{oq}|-|S_{ou}\cap S_{oq}|)/(1-P_{Mq}))-H(p_{q}))}$ (4.16)
This establishes the second case in (4.11).
Since we have already assumed that the carrier states $S_{M}(M,Tu_{m})$ and $S_{cq}(cq,Tq_{m})$ of $I_{u}=\langle ou,Tu_{h},S_{ou},M,Tu_{m},S_{M}\rangle$ and $I_{q}=\langle oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq}\rangle$ are both represent the physical realizations of $S_{ou}$ and $S_{oq}$ in model M and its interface, respectively. It follows that equation (4.11) characterizes the upper bound of the model’s TVD. This completes the proof.
Unlike traditional generalization theories such as VC dimension or Rademacher complexity, Theorem 6, based on the MLT-MF framework, provides an explicit quantitative estimate for the model’s TVD. This estimate is grounded in the overlap between the test and training data $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$, as well as the entropy $H(p_{q})$ of the test data distribution over the non-overlapping region. These metrics serve as indicators of how much the testing data shares information with the model’s prior knowledge and how unpredictable the unseen portion is.
Because TVD characterizes the alignment between distributions at a fundamental level—independent of any particular model architecture or task—it is generally applicable to the analysis of model training performance. In fact, for classification tasks with 0-1 loss, the generalization error upper bound is equivalent to the TVD Rodríguez2021 . For regression tasks, the upper bound on generalization error is given by the product of TVD and the Lipschitz constant of the loss function Devroye2013 . Thus, TVD is a meaningful and informative measure for estimating model generalization error.
This shows that Theorem 6 offers a new perspective and practical guidance for data collection, model design, and training for construction. In particular, Theorem 6 reveals that as long as the problem space is contained within the model’s information space, the model’s TVD can be made arbitrarily close to zero in a statistical sense. This is a fundamental result in machine learning:
Corollary 2 (TVD for Ideal Models) If the model’s information encompasses all relevant information for the problem, then the TVD for these problems is approximately zero.
5 Experimental Results
Using the feedforward neural network model listed in the appendix, we conducted regression experiments on two bivariate root functions: $y_{1}=\sqrt{4x_{1}^{2}+3x_{2}^{2}}$ and $y_{2}=\sqrt{2x_{1}^{2}+x_{2}^{2}}$. Both the training and test datasets were generated randomly. We evaluated the models’ prediction errors under various conditions, assessed the impact and computational cost of ethical and safety mechanisms, and explored the challenge of providing multiple types of explanations for the same output. These experiments serve to illustrate and validate the definitions and theorems presented in Section 4.
5.1 Prediction Error Experiments
To evaluate the model’s predictive performance in both overlapping and non-overlapping cases between the training and test datasets, we used the following data setup: 10,000 training samples were generated with $x_{1}$ and $x_{2}$ independently and uniformly distributed over $[0.5,1.5]$, and 10,000 test samples were drawnuniformly from within a circle of radius 1.0 centered at (1.5, 1.0). The model’s learning rate was set to $\eta=0.01$.
Figure 2 illustrates the experimental conditions and outcomes:
Figure 2: Model Prediction Error Experiments and Results Under Different Conditions. (a) The blue rectangular region represents the training dataset comprising 10,000 samples, while the red circular region represents the test dataset, also with 10,000 samples. Among these, 4,951 samples overlap with the training set. (b) For each of the training steps (10, 50, 100, 200, 500, 1000, 4000), the variation in root mean square error (RMSE) is shown for 10 randomly selected test samples from the training dataset. (c) In the non-overlapping region between the test and training datasets, the RMSEs for all 5,049 test samples are compared between the conservative uniform prediction strategy and the TVD upper bound. (d) For 10 randomly selected test samples from the non-overlapping region, the changes in RMSE are shown for training steps (10, 30, 300, 600, 1000, 4000, 7000), along with comparisons to both the TVD upper bound and the RMSE from the conservative uniform prediction.
(a) The overlap rate between the test and training datasets is approximately 50%.
(b) For 10 randomly selected samples from the overlapping region, as the number of training steps increases from 10 to 4000 (i.e., as the model becomes more sufficiently trained), the root mean square error (RMSE) of the model’s predictions steadily approaches zero. This result aligns with the first case of Theorem 6, equation (4.11).
(c) For all test samples from the non-overlapping region, a conservative uniform prediction strategy was used. For almost all samples (except a few outliers), their RMSE is less than $\text{TVD}=\sqrt{((1-0.4951)\ln((10000-4951)/(1-0.4951))-2.9741)/2}\approx 0.9155$. This indicates that even in regression tasks, although TVD is not strictly equivalent to generalization error, it can still effectively capture the distributional characteristics of generalization error.
(d) For 10 randomly selected test samples from the non-overlapping region, when using the model’s prediction method, the RMSE also decreases with more training steps, reflecting a certain degree of model generalization outside the training data region. However, the rate of error reduction is significantly lower than in the overlapping region, and for some samples, the RMSE remains higher than that of the conservative uniform prediction. In fact, when training is insufficient, the overall RMSE of the model’s predictions can even exceed the TVD. Meanwhile, the RMSE of conservative predictions for these samples is much lower than the TVD, reflecting the conclusion of the second case in Theorem 6, equation (4.11).
In summary, this regression experiment using a simple feedforward neural network model demonstrates and validates the accuracy of the model TVD upper bound estimation based on the MLT-MF framework, and its potential for estimating generalization error. Since the theoretical proof of Theorem 6 is independent of specific models and tasks, it promises to be extensible to more complex scenarios.
5.2 Ethical Safety Assurance Experiment
Given the nature of the bivariate root function regression task, the model’s predicted outputs should never be negative. This requirement is analogous to Definition 10, which ensures that model outputs do not violate the ethical constraints . Therefore, using the same training dataset as in the previous prediction error experiments, we set the learning rate to and trained the model for 165 steps. For testing, we generated a dataset of 10,000 samples with drawn uniformly from within a circle of radius 0.5 centered at (-1, -1). The criterion for judging ethical safety violations was whether the model’s predicted output was negative. This setup was used to test the feasibility of Theorem 5 and to evaluate the additional computational overhead introduced by ethical safety checking.
For each trial, 5,000 test samples were randomly selected using the trial number as a random seed. Two types of tests were conducted in parallel: “Without Ethical Safety Check” and “With Ethical Safety Check.” In the “Without Ethical Safety Check” condition, model predictions were output directly and the total computation time was recorded. In the “With Ethical Safety Check” condition, each prediction was first checked for sign; if any prediction was negative, an ethical error message was output, otherwise, the model’s prediction was output. The total computation time was also recorded in this scenario. In total, 10 trials were conducted, and the results are shown in Table 1.
Table 1: Summary Table of Ethical Safety Experiment Results
Trial Computation Computation Computation Number of Number of Correct Correct
Number Time Without Time With Time Ratio Ethically Unethical Prediction Prompt
Ethical Ethical Safe Warning Rate for Rate for
Safety Check Safety Check Samples Samples Safe Unethical
(seconds) (seconds) Outputs Outputs
1 0.367 0.391 1.063 4517 483 100% 100%
2 0.365 0.378 1.036 4521 479 100% 100%
3 0.352 0.362 1.030 4548 452 100% 100%
4 0.330 0.378 1.146 4516 484 100% 100%
5 0.371 0.383 1.032 4523 477 100% 100%
6 0.327 0.380 1.160 4546 454 100% 100%
7 0.358 0.402 1.123 4545 455 100% 100%
8 0.360 0.410 1.140 4538 462 100% 100%
9 0.361 0.365 1.010 4519 481 100% 100%
10 0.355 0.424 1.192 4533 467 100% 100%
In Table 1, the “Computation Time Ratio” is calculated as the computation time with ethical safety checking divided by the computation time without it. For the simple feedforward neural network model used in these experiments, the additional computational overhead for performing ethical safety checks before model output ranged from 1.0% to 19.2%. As model size increases, this overhead is expected to decrease in proportion; however, as the number of ethical safety constraints grows, the overhead will increase accordingly. This is a key consideration for implementing ethical safety in models. In these 10 trials, more than 19% of samples resulted in unethical predictions. In every trial, both the “Correct Prediction Rate for Safe Outputs” and the “Correct Prompt Rate for Unethical Outputs” were 100%, demonstrating that by incorporating safety checks and informative outputs, we can always guarantee the ethical safety of model predictions while maximizing the number of safe outputs.
In summary, this ethical safety experiment, grounded in Definition 10 and the formal MLT-MF-based framework for ethical safety, illustrates and validates Theorem 5 regarding the feasibility of integrating ethical safety checking mechanisms and achieving the maximal number of safe outputs.
5.3 Interpretability Experiment
For our bivariate root regression task, the output should never be negative. However, in the ethical safety assurance experiment, after 165 training steps, the model produced negative predictions for more than 19% of randomly selected 5,000-sample test batches, repeated over 10 trials. According to Theorem 4, this phenomenon can be explained from multiple perspectives, each based on different information.
5.3.1 Mathematical Operation-Based Explanation
After 165 training steps, the neural network model M has the following weight matrices and bias vectors from the input layer to the hidden layer:
$W_{1}\approx\begin{bmatrix}-0.065&-0.429\\ 0.681&-0.157\\ -0.498&-0.586\end{bmatrix},\quad b_{1}\approx\begin{bmatrix}0.129\\ -0.551\\ 0.350\end{bmatrix}$ (5.1)
The weights and biases from the hidden layer to the output layer are:
$W_{2}\approx\begin{bmatrix}0.145&0.519&-0.423\\ 0.022&0.048&0.043\end{bmatrix},\quad b_{2}\approx\begin{bmatrix}0.613\\ 0.199\end{bmatrix}$ (5.2)
Therefore, for the test sample $X=[-1.3535,-1.3535]^{T}$, the model predicts:
$Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}=\text{Identity}(W_{2}\text{ReLU}(W_{1}X+b_{1})+b_{2})\approx\begin{bmatrix}-0.040\\ 0.277\end{bmatrix}$ (5.3)
Here, $y_{1}=-0.040<0$, which mathematically explains why model M outputs a negative value for input $X$.
These formulas provide the designer’s mathematical representation of the model’s computational process. They are equivalent to the formal forward propagation rules (A.21) and (A.22) in the appendix. According to the OIT perspective, these equations describe the ontological state of the model, its outputs, and the information they convey. The model, along with its input and output mechanisms, serves as the carrier of this information and realizes the computational process. Therefore, using these formulas to explain why the model produces negative outputs satisfies the requirements of Definition 9 regarding model interpretability and meets the needs of researchers who seek to understand model behavior from a mathematical standpoint.
5.3.2 Visualization-Based Explanation of the Computational Process
Depicting the model’s computational process visually can make explanations of model behavior much more intuitive. To this end, for model M after 165 training steps, we illustrate its complete structure—covering the inputs, internal neural network calculations and data flow, and outputs—in Figure 3.
Figure 3: Schematic diagram of the computational process for model M after 165 training steps, showing its handling of the input sample $X=[-1.3535,-1.3535]^{T}$ and output $Y\approx[-0.040,0.277]^{T}$.
Compared with equations (5.1)(5.3), Figure 3 not only comprehensively reflects the input, weight matrices, bias vectors, neuron activation functions, and the specific data and formulas for outputs, but also highlights the interactions between the input/output devices and the model. The figure offers a more direct view of the internal data processing pipeline, making it easier to examine and verify each step of the computation. Thus, the visualization provides a more accessible and concrete explanation of how model M processes inputs and generates results.
5.3.3 Partition-Based Explanation
Figure 4: Illustration of Prediction Polarity and Impact of Training Steps. (a) Polarity distribution of predicted data around the training set region after model M has been trained for 165 steps. The blue square marks the area where the training dataset is located, centered at (1, 1) with a side length of 1. Green dots indicate samples with all-positive predictions; dark green dots indicate samples with all-positive predictions that also belong to the test set; red dots indicate samples with at least one negative prediction that belong to the test set; orange dots indicate samples with at least one negative prediction that do not belong to the test set. (b) Cumulative effect of up to 400 training steps on the prediction polarity at two points: (1, 1) and (-1.3535, -1.3535). The blue and green lines represent the predicted $y_{1}$ and $y_{2}$ values for the sample (1, 1) from the training set; the red and pink lines represent the predicted $y_{1}$ and $y_{2}$ values for the sample (-1.3535, -1.3535) from the test set.
It is evident that equations (5.1)(5.3) and Figure3 explain only the result of model M processing the specific sample $X=[-1.3535,-1.3535]^{T}$. They do not account for the phenomenon observed in the ethical safety assurance experiment, where each random selection of 5,000 samples from the test dataset yields over 19% negative predictions. To explain this, a broader statistical analysis across more data samples is needed.
Figure 4(a) provides a visual depiction of the decision boundary after 165 training steps, showing where model M predicts negative outputs. Notably, there is a boundary line approximately passing through points (-2.0, 0.4) and (-0.4, -2.0). For samples on the upper right side of this diagonal, model M consistently predicts non-negative values; for those on the lower left side, the model consistently outputs negative values. The test region in the figure, representing the test dataset, falls precisely on the lower left side of this boundary line, which explains why a random selection of 5,000 samples from the test set will always include a substantial portion that triggers negative predictions. This visual and statistical partitioning thus provides an intuitive explanation for the observed frequency of negative outputs in the ethical safety experiment.
Figure 4(a) presents an overall view of numerous inputs processed and predicted by model M. This visualization also represents an ontological state of M as it handles a wider range of input information—the real processing and prediction procedure of M is a direct result of this ontological mapping. Thus, according to Definition 9, Figure 4(a) constitutes a valid explanation for the distribution of prediction polarities by M.
5.3.4 Explanation Based on Cumulative Training Effects
The previous three explanations all focus on the behavior of M after 165 training steps. However, model training is a continuous process of cumulative improvement. Examining only the model’s instantaneous performance at a particular training step cannot fully explain the causes of abnormal outputs. To address this, we further trained M for 400 iterations on the training dataset. Figure 4(b) shows the predictions for both $(1,1)$ and (1.3535,1.3535)(-1.3535,-1.3535) after each training step.
It is clear that even for $(1,1)$ — a central sample from the training set — the predicted $y_{2}$ by M remains negative during roughly the first 50 training steps. Only after more than 50 steps does the prediction move away from negative values. For the test sample (1.3535,1.3535)(-1.3535,-1.3535), it takes more than about 180 training steps before the prediction begins to consistently avoid negative values. This cumulative effect is entirely due to the optimization of M’s parameters as training progresses.
In fact, by applying the forward propagation rules (A.21) and (A.22) from the appendix—substituting the model parameters obtained after each training step—one can, through a recursive process, provide a rigorous and step-by-step explanation for M’s predictions on specific input samples after any given amount of training.
6 Conclusion and Outlook
This paper has established the "Meta-Framework for Machine Learning Theory (MLT-MF)" and, through formal information modeling and the sextuple definition, has unified the characterization of information flow and causal mechanisms across the entire machine learning lifecycle. The core contributions are as follows:
  • Theoretical Integration: For the first time, data initialization, training optimization, and knowledge transfer are incorporated into a unified framework, overcoming the fragmented modeling of temporal causal chains in traditional research;
  • Interpretability: By explicitly representing ontological and carrier states via higher-order logical systems and mapping model decisions back to the formal semantic space through information recoverability, logical transparency is ensured and a rigorous mathematical foundation for interpretability is provided;
  • Ethical Safety: By embedding ethical constraint predicates, the compliance of model outputs can be formally verified, offering criteria and methodological guidance for the design of trustworthy AI systems;
  • Model Performance: Based on state coverage and KL divergence, an upper bound formula for TVD is derived, revealing the core principle that the “problem space must be contained within the model knowledge space,” thus providing explicit tools for quantitative analysis in data collection and model optimization.
We also validated and demonstrated the correctness and practicality of the relevant theorems based on MLT-MF through experiments with a simple feedforward neural network model.
Despite these important advances, the current framework assumes ideal (noiseless) enabling mappings and primarily addresses model learning and processing logic in static scenarios. It does not yet address information fusion and conflict resolution across ontological spaces in multimodal or multi-agent systems. Future research should deepen the following directions:
  • Extend to complex dynamic scenarios such as online learning under noisy conditions, introducing noisy ontologies and incremental coverage mechanisms;
  • Investigate information fusion and conflict resolution across ontological spaces for multimodal and multi-agent systems, constructing distributed MLT-MF frameworks;
  • Promote interdisciplinary integration of MLT-MF with fields such as quantum computing and bioinformatics, exploring formal mappings of quantum states and interpretable modeling of gene regulatory networks;
  • Design automated toolchains to support conversion from formal rules to code, lowering the engineering threshold for MLT-MF adoption.
7 Acknowledgements
During the writing and revision of this paper, I received many insightful comments from Associate Professor Wang Rui of the School of Computer Science at Shanghai Jiao Tong University. I also gained much inspiration and assistance from regular academic discussions with doctoral students Siyuan Qiu, Chun Li, Hu Xu, and Zeyan Li. I hereby express my sincere gratitude to them. Junxuan He from Beijing University of Posts and Telecommunications implemented all the experiments described in the paper according to the author’s ideas, producing complete model code and training/test datasets, providing indispensable and strong support for illustrating and validating the theorems related to MLT-MF. The author offers special thanks for his hard work, keen understanding, and efficient implementation.
8 Data and Code availability
All datasets and codes used in this work are available at https://github.com/xuhu6736/MLT-MF.
Appendix A Appendix: Model Instance Validation of MLT-MF
To facilitate validation and illustration of the practical significance of MLT-MF and related theorems, we select a very simple model and use MLT-MF to establish formal representations with WFFs for each stage from initial information to training information, learning optimization, and processing/output.
A.1 A Simple Feedforward Neural Network Model
We choose a feedforward neural network M with an input layer containing two inputs, a hidden layer containing three neurons, and an output layer containing two outputs (Figure A1). Its simplicity helps explain the essence of MLT-MF in a clear and straightforward manner.
Figure A1: Schematic diagram of feedforward neural network model M
The hidden layer uses the ReLU activation function:
$\text{ReLU}(x)=\max(0,x)$ (A.1)
The output layer uses the identity activation:
$\text{Identity}(x)=x$ (A.2)
Initial weight matrices and bias vectors are defined between layers:
$W_{1}=\begin{bmatrix}w_{11}^{1}&w_{12}^{1}\\ w_{21}^{1}&w_{22}^{1}\\ w_{31}^{1}&w_{32}^{1}\end{bmatrix},\quad b_{1}=\begin{bmatrix}b_{1}^{1}\\ b_{2}^{1}\\ b_{3}^{1}\end{bmatrix}$ (A.3)
$W_{2}=\begin{bmatrix}w_{11}^{2}&w_{12}^{2}&w_{13}^{2}\\ w_{21}^{2}&w_{22}^{2}&w_{23}^{2}\end{bmatrix},\quad b_{2}=\begin{bmatrix}b_{1}^{2}\\ b_{2}^{2}\end{bmatrix}$ (A.4)
Training or input data, output data, and training labels or target outputs are specified, with mean squared error (MSE) as the loss:
$X=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix},\quad Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix},\quad Y_{t}=\begin{bmatrix}y_{t1}\\ y_{t2}\end{bmatrix}\quad and\quad Y_{t}=\text{Target}(X)$ (A.5)
$L=\text{MSE}(Y,Y_{t})=\frac{1}{2}\|Y-Y_{t}\|^{2}=\frac{1}{2}((y_{1}-y_{t1})^{2}+(y_{2}-y_{t2})^{2})$ (A.6)
The gradients of the loss function $L$ with respect to the weight matrices and bias vectors are the primary basis for model M’s optimization during training. The four calculation formulas are as follows:
$\nabla_{W_{2}}L=(Y-Y_{t})(\text{ReLU}(W_{1}X+b_{1}))^{T}$ (A.7)
$\nabla_{b_{2}}L=(Y-Y_{t})$ (A.8)
$\nabla_{W_{1}}L=(W_{2}^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}X+b_{1}))X^{T}$ (A.9)
$\nabla_{b_{1}}L=W_{2}^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}X+b_{1})$ (A.10)
In equations (A.9) and A.10, the operator $\odot$ denotes element-wise multiplication, and ReLU’ represents the derivative of the ReLU function. Since ReLU is not differentiable at zero, we additionally define ReLU’(0) = 0. Thus, the update formulas for optimizing the weight matrices and bias vectors during M’s training are:
$W_{1}^{\text{new}}$ $=W_{1}-\eta\nabla_{W_{1}}L,\quad b_{1}^{\text{new}}=b_{1}-\eta\nabla_{b_{1}}L$ (A.11)
$W_{2}^{\text{new}}$ $=W_{2}-\eta\nabla_{W_{2}}L,\quad b_{2}^{\text{new}}=b_{2}-\eta\nabla_{b_{2}}L$ (A.12)
where $\eta$ is the learning rate of M.
A.2 Symbol Definitions
MLT-MF uses a formal system to express the operational logic of each phase throughout the model’s full lifecycle. Defining the symbols is a prerequisite for formal modeling. To this end, Tables A1A4 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
Table A1: Definition of Constants
Symbol Type Definition
$\mathbb{R}^{2}$ Space 2-dimensional real space
$\mathbb{R}^{3}$ Space 3-dimensional real space
Table A2: Definition of Variables
Symbol Type Definition
$i,j,k$ natural number index or count
$\varphi_{o_{i}}$ WFFs well-formed formulas contained in the informational ontological state of M’s initial model
$\psi_{o_{i}}$ WFFs well-formed formulas contained in the informational carrier state of M’s initial model
$\varphi_{t_{i}}$ WFFs well-formed formulas contained in the training informational ontological state
$\psi_{t_{i}}$ WFFs well-formed formulas contained in the training informational carrier state
$\varphi_{u_{i}}$ WFFs well-formed formulas contained in the informational ontological state of M’s optimized model
$\psi_{u_{i}}$ WFFs well-formed formulas contained in the informational carrier state of M’s optimized model
$\varphi_{q_{i}}$ WFFs well-formed formulas contained in the input informational ontological state of M
$\psi_{q_{i}}$ WFFs well-formed formulas contained in the input informational carrier state of M
$\varphi_{r_{i}}$ WFFs well-formed formulas contained in the output informational ontological state of M
$\psi_{r_{i}}$ WFFs well-formed formulas contained in the output informational carrier state of M
$\psi_{v_{i}}$ WFFs well-formed formulas contained in the informational ontological state of M’s re-optimized model
$\psi_{v_{i}}$ WFFs well-formed formulas contained in the informational carrier state of M’s reoptimized model
layerslayers natural number number of layers in the model
$v$ real number numerical values for weights and biases
$W_{k}^{(o)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the initial model
$w_{ij}^{(o)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the initial model
$b_{k}^{(o)}$ vector bias vector of neurons in the k-th layer in the initial model
$b_{i}^{(o)k}$ real number bias of the i-th neuron in the k-th layer in the initial model
$W_{k}^{(u)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the optimized model
$w_{ij}^{(u)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the optimized model
$b_{k}^{(u)}$ vector bias vector of neurons in the k-th layer in the optimized model
$b_{i}^{(u)k}$ real number bias of the i-th neuron in the k-th layer in the optimized model
$W_{k}^{(v)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the re-optimized model
$w_{ij}^{(v)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the re-optimized model
$b_{k}^{(v)}$ vector bias vector of neurons in the k-th layer in the re-optimized model
$b_{i}^{(v)k}$ real number bias of the i-th neuron in the k-th layer in the re-optimized model
$W_{k}^{(w)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the further re-optimized model
$w_{ij}^{(w)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the further re-optimized model
$b_{k}^{(w)}$ vector bias vector of neurons in the k-th layer in the further re-optimized model
$b_{i}^{(w)k}$ real number bias of the i-th neuron in the k-th layer in the further re-optimized model
$X$ vector input vector
$H$ vector hidden layer vector
$Y$ vector output vector
$Y_{t}$ vector output reference vector
$x_{i}$ real number the i-th element of the input vector
$h_{i}$ real number the i-th element of the hidden layer vector
$y_{i}$ real number the i-th element of the output vector
$y_{ti}$ real number the i-th element of the output reference vector
$t$ time time point of state occurrence
$\delta$ time time step
$L$ real number loss function value
$\nabla_{W_{1}}L$ real number gradient of the loss function with respect to $W_{1}$
$\nabla_{b_{1}}L$ real number gradient of the loss function with respect to $b_{1}$
$\nabla_{W_{2}}L$ real number gradient of the loss function with respect to $W_{2}$
$\nabla_{b_{2}}L$ real number gradient of the loss function with respect to $b_{2}$
$\eta$ Positive real number learning rate
Table A3: Definition of Functions
Symbol Number of Elements Definition
ReLU(x)(x) 1 rectified linear unit (ReLU) activation function
Identity(x)(x) 1 identity function
maxmax multivariate maximum function
[…]
ReLU’$(x)$ 1 derivative of ReLU(x)(x)
Target(X)(X) 1 objective function for input vector $X$
Table A4: Definition of Predicates
Symbol Definition
[…]
A.3 Formal Representations of Information at Each Stage of Model M
Following MLT-MF, we build formal expressions for the following five stages:
A.3.1 Initial information of M
According to the principles of OIT and based on MLT-MF, the model M is endowed with its initial information from the very beginning:
$I_{0}=(oo,To_{h},S_{oo},M,To_{m},S_{M})$ (A.13)
Here, the ontology $oo$ represents the designer’s conceptual blueprint; $To_{h}$ is the time at which the model is designed; and the ontological state $S_{oo}(oo,To_{h})$ is a set of well-formed formulas that describe the initial operational logic of model M:
$S_{oo}(oo,To_{h})=\{\varphi o_{1},\varphi o_{2},\varphi o_{3},\varphi o_{4},\varphi o_{5},\varphi o_{6},\varphi o_{7},\varphi o_{8},\varphi o_{9},\varphi o_{10},\varphi o_{11}\}$ (A.14)
The specific content of each formula is as follows:
  • Network Structure Definition
    $\varphi o_{1}=$ $\text{NetworkStructure}^{2}{(layers,3)}\land\text{LayerSize}^{2}(1,2)\land$ (A.15)
    $\text{LayerSize}^{2}(2,3)\land\text{LayerSize}^{2}(3,2)$
  • Weight Initialization
    $\varphi o_{2}$ $=\forall i\forall j(BelongsTo^{2}(i,\{1,2,3\})\land BelongsTo^{2}(j,\{1,2\})\rightarrow\exists v(Eq^{2}(w_{ij}^{(o)1},v)))$ (A.16)
    $\varphi o_{3}$ $=\forall i\forall j(BelongsTo^{2}(i,\{1,2\})\land BelongsTo^{2}(j,\{1,2,3\})\rightarrow\exists v(Eq^{2}(w_{ij}^{(o)2},v)))$ (A.17)
  • Bias Initialization
    $\varphi o_{4}=\forall iBelongsTo^{2}(i,\{1,2,3\})\rightarrow\exists v(Eq^{2}(b_{i}^{(o)1},v))$ (A.18)
    $\varphi o_{5}=\forall iBelongsTo^{2}(i,\{1,2\})\rightarrow\exists v(Eq^{2}(b_{i}^{(o)2},v))$ (A.19)
  • Activation Function Definition
    $\varphi o_{6}=\forall x(Eq^{2}(\text{ReLU}(x),max(0,x)))\land\forall x(Eq^{2}(\text{Identity}(x),x))$ (A.20)
  • Forward Propagation Rules
    $\varphi o_{7}=$ $\forall X\forall t\forall H(InputAt^{2}(X,t)\land BelongsTo^{2}(X,\mathbb{R}^{2})\rightarrow$
    $\exists H(BelongsTo^{2}(H,\mathbb{R}^{3})\land HiddenAt^{2}(H,t+\delta)\land$
    $\forall i(BelongsTo^{2}(i,\{1,2,3\})\rightarrow Eq^{2}(h_{i},\text{ReLU}\left(\sum_{j=1}^{2}w_{ij}^{(o)1}\times x_{j}+b_{i}^{(o)1}\right))))$ (A.21)
    $\varphi o_{8}=$ $\forall H\forall t(HiddenAt^{2}(H,t+\delta)\land BelongsTo^{2}(H,\mathbb{R}^{3})\rightarrow$
    $\exists Y(BelongsTo^{2}(Y,\mathbb{R}^{2})\land(OutputAt^{2}(Y,t+2\delta)\land$
    $\forall i(BelongsTo^{2}(i,\{1,2\})\rightarrow Eq^{2}(y_{i},\text{Identity}\left(\sum_{j=1}^{3}w_{ij}^{(o)2}\times h_{j}+b_{i}^{(o)2}\right)))))$ (A.22)
  • Backpropagation Rules
    $\varphi o_{9}$ $=\forall X\forall Y_{t}\forall t(InputAt^{3}(X,Y_{t},t)\land BelongsTo^{2}(Y_{t},\mathbb{R}^{2})\land\varphi o_{7}\land\varphi o_{8}\land$
    $Eq^{2}(L,\text{MSE}(Y,Y_{t}))\land Eq^{2}(\text{MSE}(Y,Y_{t}),\frac{1}{2}\sum_{i=1}^{2}(y_{i}-y_{ti})^{2}))$ (A.23)
    $\varphi o_{10}=$ $\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}\land\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),W_{1}^{(o)})\land$
    $\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),b_{1}^{(o)})\land\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),W_{2}^{(o)})\land$
    $\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),b_{2}^{(o)})\land$
    $Eq^{2}(\nabla_{W^{(o)}_{1}}L,((W_{2}^{(o)})^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))X^{T})\land$
    $Eq^{2}(\nabla_{b^{(o)}_{1}}L,((W_{2}^{(o)})^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))\land$
    $Eq^{2}(\nabla_{W^{(o)}_{2}}L,(Y-Y_{t})(\text{ReLU}(W_{1}^{(o)}X+b_{1}^{(o)}))^{T})\land Eq^{2}(\nabla_{b^{(o)}_{2}}L,(Y-Y_{t}))$ (A.24)
  • Learning Rules
    $\varphi o_{11}=$ $\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}\land\varphi o_{10}\land Eq^{2}(W_{1}^{(u)},W_{1}^{(o)}-\eta\nabla_{W^{(o)}_{1}}L)\land$
    $Eq^{2}(b_{1}^{(u)},b_{1}^{(o)}-\eta\nabla_{b^{(o)}_{1}}L)\land Eq^{2}(W_{2}^{(u)},W_{2}^{(o)}-\eta\nabla_{W^{(o)}_{2}}L)\land$
    $Eq^{2}(b_{2}^{(u)},b_{2}^{(o)}-\eta\nabla_{b^{(o)}_{2}}L)$ (A.25)
In equation (A.14), the carrier M refers to the initially constructed model, and $To_{m}$ denotes the time when model M is completed. Clearly, we have:
$\sup(To_{h})\leq\inf(To_{m})$ (A.26)
The state $S_{M}(M,To_{m})$ is a one-to-one enabling mapping of the state $S_{oo}(oo,To_{h})$ onto M, that is,
$S_{M}(M,To_{m})=\{\psi o_{1},\psi o_{2},\psi o_{3},\psi o_{4},\psi o_{5},\psi o_{6},\psi o_{7},\psi o_{8},\psi o_{9},\psi o_{10},\psi o_{11}\}$ (A.27)
and
$Io(\varphi o_{i})\underline{\triangleright}\psi o_{i},(1\leq i\leq 11)$ (A.28)
Here, the information enables the mapping from the initial ontological state of model M to its carrier state, that is, the initial state of M. In practice, this mapping is a complex process involving hardware implementation, software development, and hardware-software integration, inevitably requiring consumption of time and energy resources, and subject to various sources of noise. However, to simplify the discussion, we temporarily disregard these factors and explain the initial state of M using a completely ideal, noise-free, and recoverable information model. The subsequent discussions on other types of information will follow this approach and will not be repeated.
A.3.2 Training Information of M
Based on MLT-MF, the training information for model M is as follows:
$It=(ot,Tt_{h},S_{ot},Mt,Tt_{m},S_{Mt})$ (A.29)
Here, the ontology $ot$ refers to the object described by the training information, and the occurrence time $Tt_{h}$ denotes the point at which this object attains its corresponding state attributes. The ontological state $S_{ot}(ot,Tt_{h})$ represents the attributes of the object at time $Tt_{h}$, which are described by a set of well-formed formulas:
$S_{ot}(ot,Tt_{h})=\{\varphi t_{1},\varphi t_{2}\}$ (A.30)
  • Vector Definition
    $\varphi t_{1}=\forall X\forall Y_{t}(BelongsTo^{2}(X,\mathbb{R}^{2})\land BelongsTo^{2}(Y_{t},\mathbb{R}^{2}))$ (A.31)
  • Label Consistency
    $\varphi t_{2}=\forall X\forall Y_{t}Eq^{2}(Target(X),Y_{t})$ (A.32)
The carrier Mt serves as the interface device through which model M receives training data, and $Tt_{m}$ is the time at which model M is trained. The carrier state $S_{Mt}(Mt,Tt_{m})$ is then described by the set of well-formed formulas:
$It(\varphi t_{i})\underline{\triangleright}\psi t_{i}(1\leq i\leq 2)$ (A.33)
A.3.3 Optimized Information After M Has Learned from Training
Based on MLT-MF, after learning the training information $It$, model M updates its initial information $I_{0}$ to a new, optimized information state:
$I_{u}=(ou,Tu_{h},S_{ou},M,Tu_{m},S_{M})$ (A.34)
Here, the ontology $ou$ is a fusion of M and the object described by the training information; the occurrence time $Tu_{h}$ is the moment M learns from the training data; and the ontological state $S_{ou}(ou,Tu_{h})$ represents the theoretical state of M after training:
$S_{ou}(ou,Tu_{h})=\{\varphi u_{1},\varphi u_{6},\varphi u_{7},\varphi u_{8},\varphi u_{9},\varphi u_{10},\varphi u_{11}\}$ (A.35)
Since the structure and algorithms of M remain unchanged before and after training, $\varphi u_{1}$, $\varphi u_{6}$, and $\varphi u_{7}$ are exactly the same as $\varphi o_{1}$, $\varphi o_{6}$, and $\varphi o_{9}$ respectively. Moreover, as the weights and biases after training are already assigned values in the original information (A.25), there is no need to include equivalent formulas for $\varphi o_{2}$, $\varphi o_{3}$, $\varphi o_{4}$, $\varphi o_{5}$ in $S_{ou}(ou,Tu_{h})$. For the remaining formulas $\varphi u_{7}$, $\varphi u_{8}$, $\varphi u_{9}$, $\varphi u_{10}$, and $\varphi u_{11}$, it suffices to systematically replace $W_{k}^{(o)}$, $w_{ij}^{(o)k}$, $b_{k}^{(o)}$, $b_{i}^{(o)k}$, $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$ and $b_{i}^{(u)k}$ in $\varphi_{7}$, $\varphi_{8}$, $\varphi_{10}$, and $\varphi_{11}$ with the corresponding symbols $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$, $b_{i}^{(u)k}$, $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$ and $b_{i}^{(v)k}$ in $S_{ou}(ou,Tu_{h})$.
Similarly, the carrier state
$S_{M}(M,Tu_{m})=\{\psi u_{1},\psi u_{6},\psi u_{7},\psi u_{8},\psi u_{9},\psi u_{10},\psi u_{11}\}$ (A.36)
is a one-to-one enabling mapping from $S_{ou}(ou,Tu_{h})$ onto M, representing the state of M after completing its training.
A.3.4 Input Information for M
Processing input information is a fundamental requirement for any model. Based on the MLT-MF framework, the input information for M is defined as:
$I_{q}=(oq,Tq_{h},S_{oq},Mq,Tq_{m},S_{Mq})$ (A.37)
Here, $oq$ denotes the object described by the input information, and $Tq_{h}$ is the time at which the object’s state occurs. The state $S_{oq}(oq,Tq_{h})$ can take two possible forms:
  • The first scenario involves only input information being processed by M, in which case
    $S_{oq}(oq,Tq_{h})=\{\varphi q_{1}\}$ (A.38)
    $\varphi q_{1}=\forall XBelongsTo^{2}(X,\mathbb{R}^{2})$ (A.39)
  • The second scenario considers online learning, where
    $S_{oq}(oq,Tq_{h})=\{\varphi q_{1},\varphi q_{2}\}$ (A.40)
    $\varphi q_{1}=\forall X\forall Y_{t}(BelongsTo^{2}(X,\mathbb{R}^{2})\land BelongsTo^{2}(Y_{t},\mathbb{R}^{2}))$ (A.41)
    $\varphi q_{2}=\forall X\forall Y_{t}Eq^{2}(Target(X),Y_{t})$ (A.42)
The carrier $Mq$ for $I_{q}$ serves as the input device for M, which may or may not be the same as the carrier $Mt$ for the training information $It$. The state $S_{Mq}(Mq,Tq_{m})$ is a one-to-one enabling mapping from $S_{oq}(oq,Tq_{h})$ onto $Mq$.
A.3.5 Information After M Processes Input, Produces Output, and Undergoes Further Optimization
There are two scenarios for M after it processes input and produces output: one corresponds to the first case in A.3.4, where only output is produced without further optimization of the model itself; the other corresponds to the second case in A.3.4, where, in the context of online learning, the model both produces output and undergoes further self-optimization. Since the latter is more comprehensive, we will focus on this scenario below.
In this case, the output information $Ir=(or,Tr_{h},S_{or},Mr,Tr_{m},S_{Mr})$ has an ontological state defined as:
$S_{or}(or,Tr_{h})$ $=\{\varphi r_{1}\}=\{\varphi o_{7}\land\varphi o_{8}\land\text{OutputAt}^{2}(Y,Tr_{h})\}$ (A.43)
The carrier state $S_{Mr}(Mr,Tr_{m})$ is a one-to-one enabling mapping from $S_{or}(or,Tr_{h})$ to the output device of M.
On the other hand, since we are considering the scenario where M performs online learning, the model now carries further optimized information:
$Iv=(ov,Tv_{h},S_{ov},M,Tv_{m},S_{M})$ (A.44)
This structure is almost identical to the optimized information $Iu$ after M has learned from training. The ontological state of $Iv$ is:
$S_{ov}(ov,Tv_{h})=\{\varphi v_{1},\varphi v_{6},\varphi v_{7},\varphi v_{8},\varphi v_{9},\varphi v_{10},\varphi v_{11}\}$ (A.45)
Here, $\varphi v_{1}$, $\varphi v_{6}$, and $\varphi v_{9}$ are identical to $\varphi u_{1}$, $\varphi u_{6}$, and $\varphi u_{9}$ respectively. For the remaining formulas, it suffices to systematically replace the parameters $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$, $b_{i}^{(u)k}$, $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$ and $b_{i}^{(v)k}$ in $S_{ou}(ou,Tu_{h})$ with the corresponding parameters $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$, $b_{i}^{(v)k}$, $W_{1}^{(w)}$, $w_{ij}^{(w)k}$, $b_{k}^{(w)}$ and $b_{i}^{(w)k}$ to obtain the updated formulas, where $W_{k}^{(w)}$, $w_{ij}^{(w)k}$, $b_{k}^{(2)}$ and $b_{i}^{(w)k}$ are the weights and biases of M after further optimization.
Similarly, the carrier state
$S_{M}(M,Tv_{m})=\{\psi v_{1},\psi v_{6},\psi v_{7},\psi v_{8},\psi v_{9},\psi v_{10},\psi v_{11}\}$ (A.46)
is a one-to-one enabling mapping from $S_{ov}(ov,Tv_{h})$ onto M.
At this point, based on MLT-MF, we have constructed the formal well-formed formula representations for all ontological and carrier states of each component of the feedforward neural network model M, as illustrated in Figure A1. Although M is a simple model, it contains all the essential components of a neural network, and the discussion has not relied on any special treatment of M. This demonstrates that MLT-MF is capable of supporting the construction of well-formed, formal representations for a wide variety of neural networks, providing a solid foundation for detailed analysis and optimization.